### v2

# Projective_error_models.bib

Abstract.
We introduce more general notions of Clifford codes and stabilizer codes, the latter we call weak stabilizer codes. This is all formulated in the language of projective representation theory of finite groups and we give a novel description of the detectable errors for a Clifford code. We give a complete characterization of when a Clifford code is also a weak stabilizer code in the case where the considered error model is a nice error basis. We also give examples of infinite families of non-stabilizer Clifford codes as well as examples of non-Clifford weak stabilizer codes. The latter of these types of examples is a class of codes that have not been studied in the same systematic framework as Clifford codes and stabilizer codes.
This work was supported by The Norwegian Research Council (Project 345433) and The Norwegian National Security Authority (NSM)
1. Introduction
Quantum error correction is an important component of current scientific efforts towards developing practical realizations of quantum communications. In this article we undertake two central questions in quantum error correction. If we have a physical system that is susceptible to noise we want to try to embed our information into a subsystem, or a code space, that can protect from this noise. The central questions are then:
  1. (1)
    Given a code space, what is the noise that this code space can correct?
  2. (2)
    Given some noise we want to protect against, what is the largest code space that can detect this given noise?
Framing these questions in a mathematically rigorous way can be challenging and is highly dependent on how one chooses to model a physical system. The Knill-Laflamme conditions, cf. [KnillLaflamme97] gives an answer to the first question, which we will describe shortly. An answer to the second question for weak stabilizer codes will be described in Section 5.
The most standard way to model a quantum mechanical system is with a complex Hilbert space $V$. For our purposes $V$ will always be assumed to be finite dimensional. A quantum mechanical state can then be identified with a positive semi-definite operator $\sigma\in\mathcal{B}(V):=\{T\colon V\to V:T\text{ is linear}\}$ of trace $1$. Note that any vector $\xi\in V$ of norm $1$ gives rise to a state by considering the rank $1$ projection onto $\xi$, usually denoted $\left|\xi\right\rangle\left\langle\xi\right|$. States of this form are called pure states. Noise in the system can be modeled by a quantum channel $N$ i.e. $N\colon\mathcal{B}(V)\to\mathcal{B}(V)$ is a completely positive and trace preserving linear map. With this model, a code space is simply given by a subspace $W$ of $V$. Question (1) above can then be stated as: for which quantum channels $N\colon\mathcal{B}(V)\to\mathcal{B}(V)$ does there exist a quantum channel $R\colon\mathcal{B}(V)\to\mathcal{B}(V)$ such that
$R\circ N|_{\mathcal{B}(W)}=\operatorname{id}_{\mathcal{B}(W)}?$
Here $\mathcal{B}(W)$ is embedded into $\mathcal{B}(V)$ in the natural way. This question is completely answered by the Knill-Laflamme conditions, cf. [KnillLaflamme97]: if $\{K_{i}\}_{i=1}^{n}\subset\mathcal{B}(V)$ are the Kraus operators of $N$, i.e.
$N(\sigma)=\sum_{i=1}^{n}K_{i}\sigma K_{i}^{*},$
and $\sum_{i=1}^{m}K_{i}^{*}K_{i}=\operatorname{id}_{V}$, then there exists a quantum channel $R$ as above if and only if
$P_{W}K_{i}^{*}K_{j}P_{W}\in\mathbb{C}P_{W}$
for all $i,j\in\{1,2,\dots,n\}$. Here $P_{W}\in\mathcal{B}(V)$ denotes the orthogonal projection onto $W$.
We will take a slightly different approach. The Knill-Laflamme conditions tells us that it is essential to know which elements $X\in\mathcal{B}(V)$ satisfy
$P_{W}XP_{W}\in\mathbb{C}P_{W}.$
Elements in $\mathcal{B}(V)$ that satisfy this condition are called detectable. We will consider unitary operators on $V$ that arise as the image of a projectively faithful irreducible projective representation of a finite group. The main question we want to answer is: which of these unitaries are detectable? The reason for this is that a single unitary operator models a single error occurring in the system. To be explicit: if
$U\in U(V):=\{T\colon V\to V:T\text{ is linear and invertible with }T^{-1}=T^{*}\},$
then we can consider the quantum channel $N_{U,p}\colon\mathcal{B}(V)\to\mathcal{B}(V)$ defined by
$N_{U,p}(\sigma)=p\sigma+(1-p)U\sigma U^{*},$
where $p\in(0,1)$ is the probability that no error occurs. Note that this channel is correctable, in the sense that there exists a quantum channel $R_{U,p}\colon\mathcal{B}(V)\to\mathcal{B}(V)$ such that
$R_{U,p}\circ N_{U,p}|_{\mathcal{B}(W)}=\operatorname{id}_{\mathcal{B}(W)},$
if and only if the element $U$ is detectable (this follows by the fact that the Knill-Laflamme conditions are preserved under taking adjoints). If $\pi\colon G\to U(V)$ is a projective representation (we will define what this means in Section 2) where the size of $G$ is $n$, and $p$ is a probability distribution on the set $\{1,\dots,n\}$, then we can define the quantum channel $N_{\pi,p}\colon\mathcal{B}(V)\to\mathcal{B}(V)$ by setting
$N_{\pi,p}(\sigma)=\sum_{i=1}^{n}p(i)\pi(x_{i})\sigma\pi(x_{i})^{*},$
where $\{x_{1},\dots,x_{n}\}$ is an enumeration of $G$. It follows that the quantum channel $N_{\pi,p}$ is correctable if and only if $\pi(x_{i})$ is detectable for every $x_{i}\in G$.
Generally, the collection of unitaries that are detectable will not form a group. However, the above construction gives a way to construct correctable channels from groups of detectable unitaries.
This article is structured as follows: All the relevant definitions of projective error models and of detectable errors will be stated in Section 3. Our definition of a projective error model will essentially be the same as that of a nice error frame, cf. [ChienWaldron17]. However, we want to use the tools from projective representation theory to its full potential, and therefore choose to emphasize this aspect of the theory. We give the necessary preliminaries about projective representation theory in Section 2.
In Section 4 we define the three classes of codes we are interested in studying: stabilizer codes, Clifford codes, and weak stabilizer codes. We also prove some basic results about how these codes are related to each other. Our definition of a Clifford code generalizes the definition due to Knill, cf. [KnillI96].
We give a description of the detectable errors of Clifford codes in Section 5 in terms of the stabilizers of the code and the logical operators on the code, cf. Theorem 5.3. This result gives some useful tools to determine when a code is a Clifford code. We also give a criteria to test if a code is a weak stabilizer code, cf. Proposition 5.7.
In Section 6 we give a complete characterization of when a Clifford code is also a weak stabilizer code in the case where the projective error model is a nice error basis, cf. [KnillI96]. This recovers similar results by Klappenecker and Rötteler, and Nicolás, Martínez, and Grassl, cf. [KlappeneckerRötteler04, GrasslMartínezNicolás10].
The highlight of the next three sections is Proposition 8.2 which gives an infinite family of non-stabilizer Clifford codes in dimensions $2n$, with $n\geq 3$ being odd.
In Section 10 we present a general construction for product codes. This construction yields infinitely many more examples of Clifford codes that are not stabilizer codes.
In the final section, Section 11 we discuss some natural questions that one could consider for future research.
Acknowledgements
A lot of this work has been accomplished after many discussions with my supervisors Tron Omland, Erik Bédos, and Nadia Larsen, to whom I am extremely grateful. I would also like to thank Ningping Cao and Alexander Frei for useful discussions.
2. Preliminaries
2.1. Some notation for common groups
For an abstract cyclic group of $n$ elements we will use the notation $C_{n}$. Since $C_{n}$ can faithfully be represented as the $n$th roots of unity, we will also use $C_{n}$ to denote precisely this, i.e.,
$C_{n}:=\{z\in\mathbb{C}\colon z^{n}=1\}.$
The group operation in $C_{n}$ is denoted as multiplication.
We will also use the notation $\mathbb{Z}_{n}$ to denote a cyclic group of $n$ elements. In this case $\mathbb{Z}_{n}$ is always the set $\{0,1,2,\dots,n-1\}$. The group operation in $\mathbb{Z}_{n}$ is denoted as addition, and it is implicit that this is done modulo $n$.
We will use the notation $\mathbb{T}$ for the circle group, i.e.,
$\mathbb{T}:=\{z\in\mathbb{C}\colon|z|=1\}.$
If $G$ is a group we will use the notation $1_{G}$ to denote the identity element in $G$, or simply as $1$ when no confusion can arise. We will also use the notation $Z(G)$ to denote the center of the group $G$ i.e.,
$Z(G):=\{x\in G:xy=yx\text{ for all }y\in G\}.$
The order, or size, of a group $G$ will be denoted by $|G|$.
2.2. Projective representation theory: definitions and notation
We will only work with finite groups in this paper: whenever we write that $G$ is a group, we will assume that $G$ is finite. Furthermore, we will also assume that all representations of groups are unitary and on finite dimensional complex Hilbert spaces. All the statements we make about projective representations can be found in for example the book [CeccheriniSilbersteinTullio22].
A (linear) representation of a group $G$ on a nonzero Hilbert space $V$ is defined to be a group homomorphism $\lambda\colon G\to U(V)$, ($\lambda$ for linear). A projective representation of $G$ on $V$ is defined to be a function $\pi\colon G\to U(V)$, ($\pi$ for projective), such that the composition $q\circ\pi\colon G\to PU(V)$ is a group homomorphism. Here $PU(V)$ is the quotient of $U(V)$ by scalar multiples of the identity, and $q\colon U(V)\to PU(V)$ is the quotient map. We say that $\pi$ is projectively faithful if the composition $q\circ\pi$ is injective. That the composition $q\circ\pi$ is a group homomorphism is equivalent to saying that there exists a function $\sigma\colon G\times G\to\mathbb{T}$ such that
$\pi(x)\pi(y)=\sigma(x,y)\pi(xy),\text{ for all }x,y\in G.$
Remark 2.1.
Note that this definition differs slightly from the one found in [CeccheriniSilbersteinTullio22], where it is required that
$\pi(xy)=\sigma(x,y)\pi(x)\pi(y),\text{ for all }x,y\in G.$
This has the effect of conjugating the function $\sigma$ in our definition.
To make the dependence on the function $\sigma$ explicit, we will often call the the function $\pi$ for a $\sigma$-projective representation of $G$ on $V$. One easily shows that $\sigma$ satisfies the following equation for all $x,y,z\in G$:
$\sigma(x,y)\sigma(xy,z)=\sigma(x,yz)\sigma(y,z).$
This is called the 2-cocycle identity, and we denote the collection of all functions $G\times G\to\mathbb{T}$ that satisfy the 2-cocycle identity by $Z^{2}(G,\mathbb{T})$. Elements in $Z^{2}(G,\mathbb{T})$ are naturally called 2-cocycles.
We say that a 2-cocycle $\sigma\in Z^{2}(G,\mathbb{T})$ is a 2-coboundary if there exists a function $f\colon G\to\mathbb{T}$ such that $\sigma=\delta f$ where $\delta f\colon G\times G\to\mathbb{T}$ is the function defined by
$(\delta f)(x,y)=f(x)f(y)\overline{f(xy)},\text{ for all }x,y\in G.$
We denote the collection of all 2-coboundaries by $B^{2}(G,\mathbb{T})$. Note that $Z^{2}(G,\mathbb{T})$ becomes an abelian group when endowed with the pointwise product, and $B^{2}(G,\mathbb{T})$ is a subgroup of $Z^{2}(G,\mathbb{T})$.
We further say that two 2-cocycles $\sigma_{1},\sigma_{2}\in Z^{2}(G,\mathbb{T})$ are cohomologous if there exists a function $f\colon G\to\mathbb{T}$ such that $\sigma_{1}=(\delta f)\,\sigma_{2}$. This defines an equivalence relation on $Z^{2}(G,\mathbb{T})$, and we denote the collection of all equivalence classes under this equivalence for $H^{2}(G,\mathbb{T})$. Note that
$H^{2}(G,\mathbb{T})=Z^{2}(G,\mathbb{T})/B^{2}(G,\mathbb{T}),$
which makes it clear that $H^{2}(G,\mathbb{T})$ is an abelian group. For a 2-cocycle $\sigma\in Z^{2}(G,\mathbb{T})$ we will use the notation $[\sigma]$ to denote its class in $H^{2}(G,\mathbb{T})$. If the 2-cocycle $\sigma$ arises from a projective representation $\pi$ of $G$, then we will sometimes use the notation $[\sigma_{\pi}]$ to denote the class of $\sigma$ in $H^{2}(G,\mathbb{T})$. A nontrivial fact is that $H^{2}(G,\mathbb{T})$ is finite whenever $G$ is finite. We will make use of this fact in Proposition 3.3.
It is at this point worth mentioning that all of this terminology has a natural place in what is called group cohomology. We will need very few of the general tools from group cohomology, but the interested reader may consult [Brown82]. Note that the connection between projective representation theory and group cohomology goes back to the late 1940s, while projective representation theory for finite groups was developed by Schur in the early 1900s. See [Packer08] for a survey of projective representation theory where all these facts are mentioned.
If $\pi_{i}\colon G\to U(V_{i})$, $i=1,2$, are two $\sigma$-projective representations of $G$, we define an intertwiner from $\pi_{1}$ to $\pi_{2}$ to be a linear map $T\colon V_{1}\to V_{2}$ such that the diagram
[…]
commutes for all $x\in G$. We will often use the notation $T\colon\pi_{1}\to\pi_{2}$ to emphasize that $T$ is an intertwiner and not only a linear map. The collection of all intertwiners between $\pi_{1}$ and $\pi_{2}$ is denoted by $\operatorname{Hom}_{G}(\pi_{1},\pi_{2})$, and it is clear that this has the natural structure of a complex vector space. We say that $\pi_{1}$ is isomorphic to $\pi_{2}$ if there exists an intertwiner $T\in\operatorname{Hom}_{G}(\pi_{1},\pi_{2})$ that is also a bijection. In the case where $\pi_{1}$ is isomorphic to $\pi_{2}$ we will sometimes write $\pi_{1}\simeq\pi_{2}$.
We have that the collection of $\sigma$-projective representations of $G$ along with intertwiners form a category. This category can be identified with modules over $\mathbb{C}[G]^{\sigma}$, where $\mathbb{C}[G]^{\sigma}$ denotes the $\sigma$-twisted group algebra of $G$. As a complex vector space it is defined to be the collection of functions $F\colon G\to\mathbb{C}$. For $y\in G$, let $\delta_{y}\colon G\to\mathbb{C}$ be defined by
$\delta_{y}(x)=\begin{cases}1&\text{if }x=y,\\ 0&\text{if }x\neq y.\end{cases}$
Then $\{\delta_{y}\}_{y\in G}$ form a basis for $\mathbb{C}[G]^{\sigma}$. Sometimes, when it is clear from context, we will simply use the notation $y=\delta_{y}$. For two functions $F_{1},F_{2}\in\mathbb{C}[G]^{\sigma}$, we define their product by the following $\sigma$-twisted convolution product:
$(F_{1}*_{\sigma}F_{2})(x)=\sum_{y\in G}\sigma(y,y^{-1}x)F_{1}(y)F_{2}(y^{-1}x).$
One confirms, with some effort, that this turns $\mathbb{C}[G]^{\sigma}$ into an algebra. In the special case where $F_{2}=y$ we have that
$(F_{1}*_{\sigma}y)(x)=\sigma(xy,y)F_{1}(xy).$
Similarly, when $F_{1}=y$ we have that
$(y*_{\sigma}F_{2})(x)=\sigma(y,y^{-1}x)F_{2}(y^{-1}x).$
We can further endow $\mathbb{C}[G]^{\sigma}$ with an inner product defined by
(1) $\langle F_{1},F_{2}\rangle=\frac{1}{|G|}\sum_{x\in G}F_{1}(x)\overline{F_{2}(x)},\text{ for }F_{1},F_{2}\in\mathbb{C}[G]^{\sigma}.$
With this $\mathbb{C}[G]^{\sigma}$ may also be regarded as a Hilbert space. Note that this Hilbert space is usually denoted by $\ell^{2}(G)$. We will not use this notation, as the main thing we care about is the fact that left multiplication in $\mathbb{C}[G]^{\sigma}$ by the basis vectors $\{\delta_{y}\}_{y\in G}$ induces a $\sigma$-projective representation of $G$ (namely the left-regular $\sigma$-projective representation of $G$). The above inner product ensures that this projective representation is unitary. We will however not explicitly need that this projective representation is unitary, hence we choose to emphasize the algebraic structure involved by only using the notation $\mathbb{C}[G]^{\sigma}$.
If we have two 2-cocycles $\sigma_{1},\sigma_{2}\in Z^{2}(G,\mathbb{T})$ and a function $f\colon G\to\mathbb{T}$, then the linear map $L_{f}\colon\mathbb{C}[G]^{\sigma_{1}}\to\mathbb{C}[G]^{\sigma_{2}}$ defined by
$(L_{f}(F))(x)=f(x)\,F(x),$
is an algebra isomorphism if and only if $\sigma_{1}=(\delta f)\,\sigma_{2}$. In particular this means that the $\sigma$-projective representation theory of $G$ is only dependent on the class $[\sigma]\in H^{2}(G,\mathbb{T})$, which is a fact that is very useful when one is doing concrete computations as one can often find 2-cocycles with very nice properties that make computations easier.
2.3. Projective representation theory: irreducible representations
If $\pi$ is a projective representation of a group $G$ on a Hilbert space $V$, we say that a subspace $W$ of $V$ is $\pi$-invariant if $\pi(x)W\subset W$ for all $x\in G$. Then $\pi$ is said to be irreducible if the only $\pi$-invariant subspaces of $V$ are $V$ and $\{0\}$. We will often make use of the fact that $\pi$ is irreducible if and only if
$\operatorname{span}\{\pi(x)\colon x\in G\}=\mathcal{B}(V).$
Projective representations enjoy analogs of Maschke’s Theorem and Schur’s Lemma, cf. [CeccheriniSilbersteinTullio22]. Maschke’s Theorem says that any projective representation $\pi$ of $G$ is isomorphic to a direct sum of irreducible projective representations. Schur’s Lemma states that if $\pi_{1}$ and $\pi_{2}$ are irreducible projective representations of $G$, then
$\dim\operatorname{Hom}_{G}(\pi_{1},\pi_{2})=\begin{cases}1&\text{if }\pi_{1}\simeq\pi_{2},\\ 0&\text{if }\pi_{1}\not\simeq\pi_{2}.\end{cases}$
Projective representations also enjoy a character theory which can be used to make powerful arguments. We will need the following result, see [Cheng15, Proposition 2.2]. If $\pi_{1}$ and $\pi_{2}$ are $\sigma$-projective representations of $G$ (not necessarily irreducible), then
(2) $\langle\chi_{\pi_{1}},\chi_{\pi_{2}}\rangle=\dim\operatorname{Hom}_{G}(\pi_{1},\pi_{2}).$
Here the inner product is the one defined in Equation 1, and for $i=1,2$, $\chi_{\pi_{i}}\in\mathbb{C}[G]^{\sigma}$ are the functions defined by
$\chi_{\pi_{i}}(x)=\operatorname{Tr}(\pi_{i}(x)).$
2.4. Projective representation theory: induction and Clifford theory
Suppose that $H$ is a subgroup of $G$ and $\pi$ is a $\sigma$-projective representation of $G$ on $V$. We define a projective representation $\operatorname{Res}^{G}_{H}\pi$ of $H$ on $V$ by letting
$\operatorname{Res}^{G}_{H}\pi(x)=\pi(x),\text{ for }x\in H.$
We have that $\operatorname{Res}^{G}_{H}\pi$ is $\operatorname{Res}^{G}_{H}\sigma$-projective, where $\operatorname{Res}^{G}_{H}\sigma\in Z^{2}(H,\mathbb{T})$ is defined by restricting $\sigma$ to $H\times H$.
Induction is a procedure that is adjoint to restriction. Concretely: if $\theta$ is a $\operatorname{Res}^{G}_{H}\sigma$-projective representation of $H$ on $W$, we define a $\sigma$-projective representation $\operatorname{Ind}_{H}^{G}\theta$ of $G$ on $\operatorname{Ind}_{H}^{G}(W)$, where $\operatorname{Ind}_{H}^{G}(W)$ is the balanced tensor product:
$\operatorname{Ind}_{H}^{G}(W):=\mathbb{C}[G]^{\sigma}\otimes_{H}W.$
The subscript $H$ indicates that for any $F\in\mathbb{C}[G]^{\sigma}$, $\xi\in W$, and $x\in H$ we have that
$F\otimes(\theta(x)\xi)=(F*_{\sigma}x)\otimes\xi.$
If $\{e_{i}\}_{i=1}^{n}$ is a basis for $W$, this property ensures that $\{r\otimes e_{i}:r\in G/H,\ i\in\{1,\dots,n\}\}$ is a basis for $\operatorname{Ind}_{H}^{G}(W)$. Then for any $x\in G$ we define $\operatorname{Ind}_{H}^{G}\theta(x)$ by linearly extending the following identity:
$(\operatorname{Ind}_{H}^{G}\theta(x))(F\otimes\xi)=(x*_{\sigma}F)\otimes\xi,\text{ for }F\in\mathbb{C}[G]^{\sigma},\xi\in W.$
Since the $\sigma$-twisted convolution product in $\mathbb{C}[G]^{\sigma}$ is associative, $\operatorname{Ind}^{G}_{H}\theta\colon G\to U(\operatorname{Ind}^{G}_{H}(W))$ indeed defines a $\sigma$-projective representation of $G$.
Induced representations satisfy Frobenius reciprocity: there is a natural isomorphism
$\operatorname{Hom}_{G}(\operatorname{Ind}_{H}^{G}\theta,\pi)\simeq\operatorname{Hom}_{H}(\theta,\operatorname{Res}^{G}_{H}\pi)$
for any $\sigma$-projective representation $\pi$ of $G$ and any $\operatorname{Res}^{G}_{H}\sigma$-projective representation $\theta$ of $H$, cf. [CeccheriniSilbersteinTullio22].
We will need to introduce some more notation before we can state the results we need from Clifford theory. Assume now that $H$ is a normal subgroup of $G$, that $\pi$ is an irreducible $\sigma$-projective representation of $G$, and that $\theta$ is an irreducible $\operatorname{Res}^{G}_{H}\sigma$-projective representation of $H$. Define the inertia group of $\theta$ to be
$I_{G}(\theta):=\{x\in G\colon\theta^{x}\simeq\theta\},$
where $\theta^{x}$ is the $\operatorname{Res}^{G}_{H}\sigma$-projective representation of $H$ given by
$\theta^{x}(y)=\sigma(x^{-1},y)\overline{\sigma(x^{-1}yx,x^{-1})}\theta(x^{-1}yx),\text{ for }y\in H.$
Then the following version of Mackey’s Lemma holds: If $\ell:=\dim\operatorname{Hom}_{H}(\theta,\operatorname{Res}^{G}_{H}\pi)$ is nonzero, then
$\operatorname{Res}^{G}_{H}\pi\simeq\ell\bigoplus_{r\in G/I_{G}(\theta)}\theta^{r}.$
Finally, Clifford correspondence says the following: if $\ell=\dim\operatorname{Hom}_{H}(\theta,\operatorname{Res}^{G}_{H}\pi)$ is nonzero, then there exists a unique irreducible $\operatorname{Res}^{G}_{I_{G}(\theta)}\sigma$-projective representation $\rho$ of $I_{G}(\theta)$ such that $\operatorname{Hom}_{H}(\theta,\operatorname{Res}^{I_{G}(\theta)}_{H}\rho)$ is nonzero and
$\operatorname{Ind}_{I_{G}(\theta)}^{G}\rho\simeq\pi.$
Furthermore, $\dim\operatorname{Hom}_{H}(\theta,\operatorname{Res}^{I_{G}(\theta)}_{H}\rho)=\ell$, cf. [CeccheriniSilbersteinTullio22].
3. Error models and projective error models
If $V$ is a nonzero Hilbert space we want to consider any unitary operator on $V$ as a potential error that can occur in the system that is modeled by $V$. It is however useful to instead work with some finite amount of data instead of the whole group of unitaries $U(V)$. In this section we make precise what type of finite data of $U(V)$ we want to work with. We will also always assume that the Hilbert spaces we consider are nonzero.
Definition 3.1.
Let $V$ be a Hilbert space. If $G$ is a group admitting a faithful irreducible representation on $V$ and $\lambda$ is a choice of such a representation we say that the pair $(G,\lambda)$ is an error model on $V$. The class of all such pairs is denoted $\mathbf{EM}_{V}$.
Definition 3.2.
Let $V$ be a Hilbert space. If $G$ is a group admitting a projectively faithful irreducible projective representation on $V$ and $\pi$ is a choice of such a representation we say that the pair $(G,\pi)$ is a projective error model on $V$. The class of all such pairs is denoted $\mathbf{PEM}_{V}$.
A first question we would like to ask is what type of groups admit faithful irreducible representations, and what groups admit projectively faithful irreducible projective representations, and what are the connections between the two? For finite groups, the first question is essentially answered in [Gaschtz54]. For the whole story, in a bit more general setting, one can see [BekkadelaHarpe08, BekkadelaHarpe13]. Although it is useful to know that one can essentially classify these groups, we would like to briefly outline how the first class of groups is connected to the second class of groups.
Proposition 3.3.
Let $V$ be a Hilbert space. If $(G_{1},\lambda_{1})\in\mathbf{EM}_{V}$ then the quotient $Q:=G_{1}/Z(G_{1})$ admits a projectively faithful irreducible projective representation $\pi_{1}$ on $V$. Furthermore, the group homomorphism $q\circ\pi_{1}$ is uniquely determined by the representation $\lambda_{1}$.
Conversely, if $(G_{2},\pi_{2})\in\mathbf{PEM}_{V}$, then for any natural number $n$ such that the order of $[\sigma_{\pi_{2}}]\in H^{2}(G_{2},\mathbb{T})$ divides $n$, we have that there exists a central extension $E$ of $G_{2}$ by $C_{n}$ that admits a faithful irreducible representation on $V$.
Proof.
For the first statement, let $(G_{1},\lambda_{1})\in\mathbf{EM}_{V}$. We will define a projective representation of $Q$ on $V$. For each $x\in Q$ let $s_{x}\in G_{1}$ be any coset representative of $x$. For $x,y\in Q$ we have that there exists a unique element $\sigma(x,y)\in Z(G_{1})$ such that $s_{x}s_{y}=\sigma(x,y)s_{xy}$. Since $\lambda_{1}$ is an irreducible representation of $G_{1}$ we have by Schur’s Lemma that $\lambda_{1}(\sigma(x,y))$ is a scalar multiple of $1_{U(V)}$. By abuse of notation we also denote this scalar by $\sigma(x,y)$. We now define a function $\pi_{1}\colon Q\to U(V)$ by the equation
$\pi_{1}(x):=\lambda_{1}(s_{x}).$
Then $\pi_{1}$ is a $\sigma$-projective representation of $Q$. Note that by construction we have that the diagram
[…]
commutes. By surjectivity of the quotient map $p\colon G_{1}\to Q$ we get that $q\circ\pi_{1}$ is the unique group homomorphism making this diagram commute.
It remains to check that $\pi_{1}$ is projectively faithful: Let $x\in\ker{(q\circ\pi_{1})}$. By assumption we have that
$\lambda_{1}(s_{x})=\pi_{1}(x)=z1_{U(V)}$
for some $z\in\mathbb{T}$. Hence, for any $g\in G_{1}$ we have that
$\lambda_{1}(s_{x}g)=\lambda_{1}(gs_{x}).$
Since $\lambda_{1}$ is injective ($\lambda_{1}$ is a faithful representation of $G_{1}$ on $V$) we have that $s_{x}g=gs_{x}$ for every $g\in G_{1}$. Thus, $s_{x}\in Z(G_{1})$, meaning that $x=1_{Q}$. Hence, $q\circ\pi_{1}$ is injective.
For the second statement, let now $(G_{2},\pi_{2})\in\mathbf{PEM}_{V}$, and suppose that $\pi_{2}$ is a $\sigma$-projective representation of $G_{2}$ on $V$. Let $n$ be a natural number such that the order of $[\sigma]\in H^{2}(G_{2},\mathbb{T})$ divides $n$. The following short exact sequence,
[…]
induces a long exact sequence in group cohomology with the following segment,
[…]
where $n$ here denotes the map that is multiplication by $n$. The fact that the order of $[\sigma]$ divides $n$ is precisely the fact that $[\sigma]\in\ker{n}$, and since this sequence is exact, this means that there exists an element $[\sigma^{\prime}]\in H^{2}(G,C_{n})$ such that $i^{*}([\sigma^{\prime}])=[\sigma]$. Fix a representative $\sigma^{\prime}\colon G_{2}\times G_{2}\to C_{n}$ of $[\sigma^{\prime}]$ and choose a function $f\colon G_{2}\to\mathbb{T}$ such that $\sigma^{\prime}=(\delta f)\,\sigma$. Then we have that
$f(x)\pi_{2}(x)\,f(y)\pi_{2}(y)=\sigma^{\prime}(x,y)f(xy)\pi_{2}(xy)\text{ for all }x,y\in G_{2}.$
Let $E$ denote the central extension of $G_{2}$ by $C_{n}$ corresponding to the element $[\sigma^{\prime}]\in H^{2}(G_{2},C_{n})$. Then define a representation $\lambda_{2}\colon E\to U(V)$ by
$\lambda_{2}(z,x)=zf(x)\pi_{2}(x)$
for $z\in C_{n}$ and $x\in G_{2}$. This is easily seen to define a faithful irreducible representation of $E$ on $V$, completing the proof. ∎
Given an error model $(G,\lambda)\in\mathbf{EM}_{V}$ and a code space $W\subset V$, one of the most important quantities to study is the collection of detectable errors in $G$ over $W$. This will be the elements of $G$ that satisfy the Knill-Laflamme conditions, cf. [KnillLaflamme97]. To be precise we make the following definitions.
Definition 3.4.
Let $V$ be a Hilbert space and $W\subset V$ be a subspace. If $(G_{1},\lambda_{1})\in\mathbf{EM}_{V}$ we define the set of detectable errors in $G_{1}$ over $W$ to be the set
$D_{(G_{1},\lambda_{1})}(W):=\{x\in G_{1}\colon P_{W}\lambda_{1}(x)P_{W}\in\mathbb{C}P_{W}\},$
where $P_{W}$ denotes the orthogonal projection onto $W$. If $(G_{2},\pi_{2})\in\mathbf{PEM}_{V}$ we analogously define the set
$D_{(G_{2},\pi_{2})}(W)=\{x\in G_{2}\colon P_{W}\pi_{2}(x)P_{W}\in\mathbb{C}P_{W}\}.$
With this we can have a precise formulation of the questions we want to answer. Given an error model $(G,\lambda)\in\mathbf{EM}_{V}$ we want to answer:
  1. (1)
    If $W\subset V$ is a subspace, can we gain an alternative description of $D_{(G,\lambda)}(W)$?
  2. (2)
    If $D\subset G$ is some subset, what is the largest (in terms of dimension) subspace $W$ of $V$ such that $D\subset D_{(G,\lambda)}(W)$, and can we describe this subspace $W$?
Of course we are also interested in the analogous questions for projective error models, and in some sense, we are only interested in these questions for projective error models:
Proposition 3.5.
Let $V$ be a Hilbert space and $W\subset V$ be a subspace. If $(G,\lambda)\in\mathbf{EM}_{V}$ then
$D_{(G,\lambda)}(W)=\{x\in G\colon p(x)\in D_{(G/Z(G),\pi)}(W)\},$
where $p\colon G\to G/Z(G)$ denotes the quotient map, and $\pi\colon G/Z(G)\to U(V)$ is any projectively faithful irreducible projective representation such that $q\circ\pi\circ p=q\circ\lambda$.
Proof.
Let $(G,\lambda)\in\mathbf{EM}_{V}$ and $\pi\colon G/Z(G)\to U(V)$ be any projectively faithful irreducible projective representation such that $q\circ\pi\circ p=q\circ\lambda$. Note that the existence of such a projective representation $\pi$ is guaranteed by Proposition 3.3. Then for any $x\in G$ we have that there exists an element $f(x)\in\mathbb{T}$ such that
$\pi(p(x))=f(x)\lambda(x).$
From this it is clear that $x\in D_{(G,\lambda)}(W)$ if and only if $p(x)\in D_{(G/Z(G),\pi)}(W)$, completing the proof. ∎
Note that this also shows that for any projective error model $(G,\pi)\in\mathbf{PEM}_{V}$, and any subspace $W\subset V$, the collection of detectable errors $D_{(G,\pi)}(W)$ is only dependent on the group homomorphism $q\circ\pi\colon G\to PU(V)$, and not on the projective representation $\pi$ of $G$ on $V$. One could therefore consider defining a projective error model on $V$ to be a pair $(G,\pi^{\prime})$ where $G$ is a finite group and $\pi^{\prime}\colon G\to PU(V)$ is a group homomorphism satisfying some niceness properties analogous to irreducibility and faithfulness. The problem with this is that one does not have a good notion of a representation theory for such group homomorphisms. The reason for this is that there might be inequivalent projective representations $\pi_{1},\pi_{2}\colon G\to U(V)$ such that
$q\circ\pi_{1}=q\circ\pi_{2}=\pi^{\prime}.$
This happens for example for the group $D_{4}$, the dihedral group of $8$ elements, which has precisely two irreducible projective representations of degree 2 (with a nontrivial 2-cocycle). Both of these representations give the same group homomorphism when precomposed with the quotient map as above. A way to see that this is the case is by looking at the character table for $D_{4}=\langle a,b\ |\ a^{4}=b^{2}=1,\ bab=a^{-1}\rangle$:
$\begin{array}[]{|c||r||r|r||r||r|r||r|r|}\hline\cr\rm Elements&1&a&a^{3}&a^{2}&b&a^{2}b&ab&a^{3}b\cr\hline\cr\hline\cr\rho_{1}&1&1&1&1&1&1&1&1\cr\hline\cr\rho_{2}&1&1&1&1&-1&-1&-1&-1\cr\hline\cr\rho_{3}&1&-1&-1&1&1&1&-1&-1\cr\hline\cr\rho_{4}&1&-1&-1&1&-1&-1&1&1\cr\hline\cr\rho_{5}&2&0&0&-2&0&0&0&0\cr\hline\cr\hline\cr\chi_{1}&2&1+i&1-i&0&0&0&0&0\cr\hline\cr\chi_{2}&2&-1-i&-1+i&0&0&0&0&0\cr\hline\cr\end{array}$
Here we have grouped the elements of $D_{4}$ according to their conjugacy classes, the $\rho_{i}$’s are all the linear characters of $D_{4}$, and $\chi_{1}$ and $\chi_{2}$ are the two projective characters of $D_{4}$, where $\chi_{1}$ is the character associated to the projective representation described in Proposition 7.2. Let $\pi_{1}\colon D_{4}\to U(\mathbb{C}^{2})$ be a projective representation whose character is $\chi_{1}$. Since the degree of $\rho_{3}$ above is $1$, we have that $\rho_{3}$ is in fact a group homomorphism, hence the projective representation $\pi_{2}\colon D_{4}\to U(\mathbb{C}^{2})$ defined by
$\pi_{2}(x)=\rho_{3}(x)\pi_{1}(x)$
has the same 2-cocycle associated with it as $\pi_{1}$. Clearly we have that $q\circ\pi_{1}=q\circ\pi_{2}$ since $\rho_{3}(x)\in\mathbb{T}$ for every $x\in D_{4}$. But the character of $\pi_{2}$ is necessarily given by $\rho_{3}\,\chi_{1}=\chi_{2}$, hence $\pi_{1}$ and $\pi_{2}$ are not isomorphic.
Next we present the most well-known example of an error model.
Example 3.6.
Consider the Pauli group $P_{n}$ on $n$ qubits given as a subgroup of $U((\mathbb{C}^{2})^{\otimes n})$. It is an error model, in the sense of Definition 3.1, by considering the pair $(P_{n},i)$, where $i\colon P_{n}\to U((\mathbb{C}^{2})^{\otimes n})$ is the inclusion homomorphism. Physically, the Pauli group is an error model that models errors occurring independently on different qubits, this is typical of product error models which we will get back to.
A projective error model associated to the Pauli error model is given by $(G,\pi)$ where $G=(\mathbb{Z}_{2}\times\mathbb{Z}_{2})^{n}$ and $\pi$ is the $n$-fold product representation of the only projectively faithful irreducible projective representation of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. Explicitly, if
$X=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},\ Z=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix},$
we define $\pi\colon(\mathbb{Z}_{2}\times\mathbb{Z}_{2})^{n}\to U((\mathbb{C}^{2})^{\otimes n})$ by
$\pi(a_{1},b_{1},\dots,a_{n},b_{n})=X^{a_{1}}Z^{b_{1}}\otimes\cdots\otimes X^{a_{n}}Z^{b_{n}}.$
The Pauli group is a very important example of an error model. The first versions of more general error models are due to Knill, cf. [KnillI96, KnillII96] where the term nice error basis is introduced. In our setup, a nice error basis is a projective error model $(G,\pi)\in\mathbf{PEM}_{V}$ such that $|G|=(\dim V)^{2}$. In the literature, groups that admit an irreducible projective representation with this property are sometimes called groups of central type. The reader should be warned that this term is sometimes also used to refer to a group $E$ such that the quotient $E/Z(E)$ admits an irreducible projective representation on a Hilbert space $V$ with the property that $|E/Z(E)|=(\dim V)^{2}$.
In [ChienWaldron17] the notion of a nice error frame is introduced. This is precisely the data of a projective error model as above. For a more thorough account of nice error frames, one can consult the book [Waldron18].
4. Stabilizer codes, Clifford codes, and weak stabilizer codes: definitions
Given an error model $(G,\lambda)\in\mathbf{EM}_{V}$, the standard way of producing error correcting codes for this model is by using some additional algebraic data, usually in the form of a normal subgroup $N$ of $G$ and an irreducible representation of $N$. The most well-known codes that arise in this way are the stabilizer codes, which were first studied in the concrete case where the error model is the Pauli group, cf. [Gottesman97, Calderbank97], where the normal subgroup $N$ was assumed to be abelian. Generalizations of this were quickly developed, namely Clifford codes, cf. [KnillI96, KlappeneckerRöttelerI02, KlappeneckerRöttelerII02], where the error models they considered were nice error bases and any normal subgroup was considered, not just abelian ones.
We will define some codes that generalize these existing codes, and we will study how they are related to each other. All our definitions will be made using projective error models. Note that completely analogous definitions can be made with error models in place of projective error models.
Definition 4.1.
Let $V$ be a Hilbert space and $(G,\pi)\in\mathbf{PEM}_{V}$. A nonzero subspace $W\subset V$ is a stabilizer code with respect to the pair $(G,\pi)$ if there exists a normal subgroup $N$ of $G$ and a function $f\colon N\to\mathbb{T}$ such that
$W=V^{St}(G,\pi,N,f):=\{\xi\in V\colon\pi(x)\xi=f(x)\xi\text{ for all }x\in N\}.$
It will often be clear from context which projective error model we are considering. In this case we will simply refer to $W$ as a stabilizer code.
Note that the requirement of $N$ being a normal subgroup of $G$ is what is usually considered in the literature, see for example [KlappeneckerRöttelerII02]. It is possible to consider the same setup with general subgroups of $G$, and we therefore make the following definition.
Definition 4.2.
Let $V$ be a Hilbert space and $(G,\pi)\in\mathbf{PEM}_{V}$. A nonzero subspace $W\subset V$ is a weak stabilizer code with respect to the pair $(G,\pi)$ if there exists a subgroup $H$ of $G$ and a function $f\colon H\to\mathbb{T}$ such that
$W=V^{St}(G,\pi,H,f)=\{\xi\in V\colon\pi(x)\xi=f(x)\xi\text{ for all }x\in H\}.$
We will often only refer to $W$ as a weak stabilizer code.
In Section 9 we present a family of codes that are weak stabilizer codes but not stabilizer codes. Hence, this weaker definition does capture more examples of codes than the definition of a stabilizer code.
It is clear that any stabilizer code is automatically a weak stabilizer code. We can also characterize when the space $V^{St}(G,\pi,H,f)$ is nonzero. First we prove some auxiliary results.
Proposition 4.3.
Let $V$ be a Hilbert space, $(G,\pi)\in\mathbf{PEM}_{V}$, $H$ a subgroup of $G$, and $f\colon H\to\mathbb{T}$ a function. If the subspace $V^{St}(G,\pi,H,f)$ is nonzero, then $f$ is a projective representation of $H$ on $\mathbb{C}$. If $\pi$ is $\sigma$-projective, then $f$ is $\operatorname{Res}^{G}_{H}\sigma$-projective.
Proof.
Since $V^{St}(G,\pi,H,f)$ is nonzero, there exists a nonzero $\xi\in V^{St}(G,\pi,H,f)$, and for every $x,y\in H$ we have that
$f(x)f(y)\xi=\pi(x)\pi(y)\xi=\sigma(x,y)\pi(xy)\xi=\sigma(x,y)f(xy)\xi.$
Hence, $f(x)f(y)=\sigma(x,y)f(xy)$ for all $x,y\in H$, completing the proof. ∎
Corollary 4.4.
Let $V$ be a Hilbert space, $(G,\pi)\in\mathbf{PEM}_{V}$, $H$ a subgroup of $G$, and $f\colon H\to\mathbb{T}$ a function. Suppose that $\pi$ is $\sigma$-projective. If the subspace $V^{St}(G,\pi,H,f)$ is nonzero, then $\operatorname{Res}^{G}_{H}\sigma$ is a coboundary.
Proof.
By Proposition 4.3 we have that $\operatorname{Res}^{G}_{H}\sigma=\delta f$, completing the proof. ∎
Theorem 4.5.
Let $V$ be a Hilbert space, $(G,\pi)\in\mathbf{PEM}_{V}$ with $\pi$ being $\sigma$-projective, and $H$ a subgroup of $G$. Consider the following statements:
  1. (1)
    There exists a function $f\colon H\to\mathbb{T}$ such that the subspace $V^{St}(G,\pi,H,f)$ is nonzero.
  2. (2)
    There exists a function $f\colon H\to\mathbb{T}$ such that $\operatorname{Res}^{G}_{H}\sigma=\delta f$ and the vector space of intertwiners $\operatorname{Hom}_{H}(f,\operatorname{Res}^{G}_{H}\pi)$ is nonzero.
  3. (3)
    The subgroup $H$ is abelian and $\operatorname{Res}^{G}_{H}\sigma$ is a coboundary.
We have that $(2)$ is equivalent to $(1)$, and that $(3)$ implies $(2)$. Furthermore, if $H$ is a normal subgroup of $G$, then $(2)$ implies $(3)$.
Proof.
That (1) implies (2) is Proposition 4.3 and Corollary 4.4. To show that (2) implies (1) let $T\in\operatorname{Hom}_{H}(f,\operatorname{Res}^{G}_{H}\pi)$ be a nonzero intertwiner. Then $\xi:=T(1)\in V$ is nonzero, and since $T$ is an intertwiner we have that
$\pi(x)\xi=\pi(x)T(1)=T(f(x)1)=f(x)T(1)=f(x)\xi,\text{ for all }x\in H.$
Hence, $\xi\in V^{St}(G,\pi,H,f)$, showing that this space is nonzero.
To show that (3) implies (2), write $\operatorname{Res}^{G}_{H}\sigma=\delta f^{\prime}$ for some function $f^{\prime}\colon H\to\mathbb{T}$. Then the map $(\overline{f^{\prime}}\,\operatorname{Res}^{G}_{H}\pi)\colon H\to U(V)$ defined by
$(\overline{f^{\prime}}\,\operatorname{Res}^{G}_{H})(x)=\overline{f^{\prime}(x)}\pi(x),$
is a group homomorphism, or rather a representation of $H$ on $V$. Since $H$ is abelian we can write $\overline{f^{\prime}}\,\operatorname{Res}^{G}_{H}\pi$ as a direct sum of 1-dimensional representations. Let $\chi\colon H\to\mathbb{T}$ be one of these representations. By choice of $\chi$ we have that $\operatorname{Hom}_{H}(\chi,\overline{f^{\prime}}\,\operatorname{Res}^{G}_{H}\pi)$ is nonzero, hence the vector space of ($\operatorname{Res}^{G}_{H}\sigma$-projective) intertwiners $\operatorname{Hom}_{H}(f^{\prime}\chi,\operatorname{Res}^{G}_{H}\pi)$ is nonzero. Define $f=f^{\prime}\chi$. Then we have that
$\operatorname{Res}^{G}_{H}\sigma=\delta f^{\prime}=(\delta\chi)\,(\delta f^{\prime})=\delta(\chi f^{\prime})=\delta f,$
showing that (3) indeed implies (2).
Now assume that $H$ is a normal subgroup of $G$. To show that (2) implies (3), note that we assume that the vector space of intertwiners $\operatorname{Hom}_{H}(f,\operatorname{Res}^{G}_{H}\pi)$ is nonzero. Hence, by Mackey’s Lemma we have that
$\operatorname{Res}^{G}_{H}\pi\simeq\ell\bigoplus_{r\in G/I_{G}(f)}f^{r},$
where $\ell=\dim\operatorname{Hom}_{H}(f,\operatorname{Res}^{G}_{H}\pi)$. Since all the $\operatorname{Res}^{G}_{H}\sigma$-projective representations $f^{r}$ are 1-dimensional we can make the following computation for each $x,y\in H$,
$\sigma(x,y)\pi(xy)$ $=\ell\bigoplus_{r\in G/I_{G}(f)}\sigma(x,y)f^{r}(xy)$
$=\ell\bigoplus_{r\in G/I_{G}(f)}f^{r}(x)f^{r}(y)$
$=\ell\bigoplus_{r\in G/I_{G}(f)}f^{r}(y)f^{r}(x)$
$=\ell\bigoplus_{r\in G/I_{G}(f)}\sigma(y,x)f^{r}(yx)$
$=\sigma(y,x)\pi(yx).$
Since $\pi$ is a projectively faithful $\sigma$-projective representation of $G$ we get that $xy=yx$ for all $x,y\in H$. Hence, $H$ is abelian, completing the proof. ∎
Corollary 4.6.
Let $V$ be a Hilbert space. Suppose that $(G,\pi)\in\mathbf{PEM}_{V}$ is a projective error model with $G$ being abelian and $\pi$ being $\sigma$-projective. Suppose that $H$ is a subgroup of $G$. Then there exists a function $f\colon H\to\mathbb{T}$ such that $V^{St}(G,\pi,H,f)$ is nonzero if and only if $\operatorname{Res}^{G}_{H}\sigma$ is a coboundary.
Proof.
Since any subgroup of $G$ is automatically normal and abelian, we have that there exists a function $f\colon H\to\mathbb{T}$ such that $V^{St}(G,\pi,H,f)$ is nonzero if and only if $\operatorname{Res}^{G}_{H}\sigma$ is a coboundary by Theorem 4.5. ∎
Proposition 4.7.
Let $V$ be a Hilbert space and $(G,\pi)\in\mathbf{PEM}_{V}$. If $H$ is a subgroup of $G$ and $f\colon H\to\mathbb{T}$ is a function such that $V^{St}(G,\pi,H,f)$ is nonzero, then
$\dim V^{St}(G,\pi,H,f)=\dim\operatorname{Hom}_{H}(f,\operatorname{Res}^{G}_{H}\pi).$
Proof.
By Proposition 4.3 $f$ is a projective representation of $H$, hence we may consider the vector space of intertwiners $\operatorname{Hom}_{H}(f,\operatorname{Res}^{G}_{H}\pi)$, and by Theorem 4.5, this space is nonzero. We claim that the map $\operatorname{Hom}_{H}(f,\operatorname{Res}^{G}_{H}\pi)\to V^{St}(G,\pi,H,f)$ defined by $T\mapsto T(1)$ is a linear isomorphism.
It is clear that this map is linear and injective, and by the same argument as in the proof of Theorem 4.5, we have that $T(1)$ is indeed an element of $V^{St}(G,\pi,H,f)$. To see that it is surjective, suppose that $\xi\in V^{St}(G,\pi,H,f)$. Define $T_{\xi}\colon\mathbb{C}\to V^{St}(G,\pi,H,f)$ by
$T_{\xi}(z)=z\xi,\text{ for }z\in\mathbb{C}.$
Then for any $x\in H$,
$T_{\xi}(f(x))=f(x)\xi=\pi(x)\xi=\pi(x)T_{\xi}(1).$
Hence, $T_{\xi}\in\operatorname{Hom}_{H}(f,\operatorname{Res}^{G}_{H}\pi)$, completing the proof. ∎
Note that stabilizer codes and weak stabilizer codes give a way to construct the largest code space that is stabilized by a subgroup of $G$. Clifford codes, which we define next, are qualitatively different as they are constructed from a subgroup of $G$ that will correspond to the logical operations on the code.
Definition 4.8.
Let $V$ be a Hilbert space and $(G,\pi)\in\mathbf{PEM}_{V}$. A subspace $W\subset V$ is a Clifford code with respect to the pair $(G,\pi)$ if there exists a subgroup $I$ of $G$ such that $W$ is:
  • $(\operatorname{Res}^{G}_{I}\pi)$-.invariant,
  • the projective representation $(\operatorname{Res}^{G}_{I}\pi)|_{W}$ is irreducible, and
  • $\operatorname{Ind}^{G}_{I}((\operatorname{Res}^{G}_{I}\pi)|_{W})\simeq\pi$.
We will often refer to $W$ as a Clifford code.
Proposition 4.9.
Let $V$ be a Hilbert space and $(G,\pi)\in\mathbf{PEM}_{V}$. If $I$ is a subgroup of $G$ and $\rho\colon I\to U(V_{\rho})$ is an irreducible projective representation such that $\operatorname{Ind}^{G}_{I}\rho\simeq\pi$, then there exists a unique subspace $V^{Cl}(G,\pi,I,\rho)$ of $V$ with the property that $\operatorname{Res}^{G}_{I}\pi$ restricted to $V^{Cl}(G,\pi,I,\rho)$ is an irreducible projective representation of $I$ and $(\operatorname{Res}^{G}_{I}\pi)|_{V^{Cl}(G,\pi,I,\rho)}\simeq\rho$.
Proof.
Let $T\in\operatorname{Hom}_{I}(\rho,\operatorname{Res}^{G}_{I}\pi)$. Since $\ker T$ is an $I$-invariant subspace of $V_{\rho}$, we have that $T$ is either injective or the 00 intertwiner. By Frobenius reciprocity and Schur’s Lemma we have that
$\dim\operatorname{Hom}_{I}(\rho,\operatorname{Res}^{G}_{I}\pi)=\dim\operatorname{Hom}_{G}(\operatorname{Ind}^{G}_{I}\rho,\pi)=1.$
Hence, there exists a nonzero intertwiner $T\colon V_{\rho}\to V$, and this intertwiner is unique up to scaling. Define $V^{Cl}(G,\pi,I,\rho):=T(V_{\rho})$. Since $T$ is unique up to scaling, we have that $T(V_{\rho})=T^{\prime}(V_{\rho})$ for any other nonzero intertwiner $T^{\prime}$, making the subspace $V^{Cl}(G,\pi,I,\rho)$ independent of choice of intertwiner.
Lastly, it is clear that $\rho\simeq(\operatorname{Res}^{G}_{I}\pi)|_{V^{Cl}(G,\pi,I,\rho)}$, namely $T\colon V_{\rho}\to V^{Cl}(G,\pi,I,\rho)$ is an intertwiner displaying this isomorphism. This completes the proof. ∎
Proposition 4.10.
Let $V$ be a Hilbert space and $(G,\pi)\in\mathbf{PEM}_{V}$. A subspace $W\subset V$ is a Clifford code if and only if there exist a subgroup $I$ of $G$ and an irreducible projective representation $\rho\colon I\to U(V_{\rho})$ with $\operatorname{Ind}_{I}^{G}\rho\simeq\pi$ and $W=V^{Cl}(G,\pi,I,\rho)$.
Proof.
It is clear from Proposition 4.9 that if $W=V^{Cl}(G,\pi,I,\rho)$, then $W$ is a Clifford code as in Definition 4.8. On the other hand, if $W$ is a Clifford code, then there exists a subgroup $I$ of $G$ such that $W$ is $(\operatorname{Res}^{G}_{I}\pi)$-invariant and irreducible, with $\operatorname{Ind}^{G}_{I}((\operatorname{Res}^{G}_{I}\pi)|_{W})\simeq\pi$. We would like to show that the inclusion map $i\colon W\to V$ is an intertwiner in $\operatorname{Hom}_{I}((\operatorname{Res}^{G}_{I}\pi)|_{W},\operatorname{Res}^{G}_{I}\pi)$. In that case,
$W=i(W)=V^{Cl}(G,\pi,I,(\operatorname{Res}^{G}_{I}\pi)|_{W})$
by the definition of the space $V^{Cl}(G,\pi,I,(\operatorname{Res}^{G}_{I}\pi)|_{W})$. But $W$ is assumed to be $(\operatorname{Res}^{G}_{I}\pi)$-invariant, hence the inclusion is an intertwiner, completing the proof. ∎
Remark 4.11.
Note that our definition of a Clifford code is slightly more general than that found in [KnillI96, KlappeneckerRöttelerI02, KlappeneckerRöttelerII02]. Their initial data to construct the space $V^{Cl}(G,\pi,I,\rho)$ is a normal subgroup $N$ of $G$ and an irreducible projective representation $\theta$ of $N$ such that the vector space of intertwiners $\operatorname{Hom}_{N}(\theta,\operatorname{Res}^{G}_{N}\pi)$ is nonzero. The subgroup $I$ of $G$ is then chosen to be the inertia group of $\theta$, $I_{G}(\theta)$. Then they require that $\rho$ is the unique irreducible projective representation of $I_{G}(\theta)$ such that $\operatorname{Hom}_{N}(\theta,\operatorname{Res}^{I}_{N}\rho)$ is nonzero and $\operatorname{Ind}^{G}_{I}\rho\simeq\pi$. Such a projective representation is guaranteed to exists by Clifford correspondence.
To end this section we show that every stabilizer code is a Clifford code.
Proposition 4.12.
Let $V$ be a Hilbert space, $(G,\pi)\in\mathbf{PEM}_{V}$, $N$ be a normal subgroup of $G$, and $f\colon N\to\mathbb{T}$ be a function. If $W=V^{St}(G,\pi,N,f)$ is a stabilizer code, then $W$ is a Clifford code.
Proof.
Since $W$ is a stabilizer code, we have by definition that $V^{St}(G,\pi,N,f)$ is nonzero. Thus, by the equivalence of (1) and (2) in Theorem 4.5 we have that the vector space of intertwiners $\operatorname{Hom}_{H}(f,\operatorname{Res}^{G}_{N}\pi)$ is nonzero. Let $I=I_{G}(f)$. Then by Clifford correspondence there exist a unique irreducible projective representation $\rho\colon I\to U(V_{\rho})$ such that $\operatorname{Ind}^{G}_{I}\rho\simeq\pi$ and the vector space of intertwiners $\operatorname{Hom}_{N}(f,\operatorname{Res}^{I}_{N}\rho)$ is nonzero. Let $W^{\prime}=V^{Cl}(G,\pi,I,\rho)$. By Proposition 4.10 we have that
$\rho\simeq(\operatorname{Res}^{G}_{I}\pi)|_{W^{\prime}}.$
Combining this with Mackey’s Lemma we have that
$(\operatorname{Res}^{G}_{N}\pi)|_{W^{\prime}}\simeq\operatorname{Res}^{I}_{N}\rho\simeq\ell f,$
where $\ell=\dim\operatorname{Hom}_{N}(f,\operatorname{Res}^{I}_{N}\rho)$. Hence, for every $\xi\in W^{\prime}$ we obtain
$\pi(x)\xi=f(x)\xi,\text{ for all }x\in N.$
Hence, $W^{\prime}\subset W$. By Clifford correspondence we also have that $\dim\operatorname{Hom}_{N}(f,\operatorname{Res}^{G}_{N}\pi)=\ell$, hence $\dim W^{\prime}=\dim W$. Thus, $W^{\prime}=W$ completing the proof. ∎
To summarize, we have defined three different types of codes: stabilizer codes, weak stabilizer codes, and Clifford codes. They are related in the following heuristic:
$\{\text{Clifford codes}\}\supset\{\text{Stabilizer codes}\}\subset\{\text{Weakstabilizer codes}\}.$
These inclusions are strict in general, which we will see in Section 8 and Section 9. Some interesting questions are
  1. (1)
    Do we have
[…]
  2. (2)
    Are there good criteria to determine when a Clifford code is a stabilizer code?
  3. (3)
    And similarly, are there good criteria to determine when a weak stabilizer code is a stabilizer code?
Questions (1) and (3) appear to be new. Question (2) is answered in [KlappeneckerRötteler04] and [GrasslMartínezNicolás10] for groups of central type, we will get back to this question in Section 6.
5. Clifford codes: error detection
In this section we want to study the collection of detectable errors for Clifford codes. By Proposition 4.12 this will also give a description of the detectable errors for a stabilizer code. We first make a couple of definitions to establish some notation.
Definition 5.1.
Let $V$ be a Hilbert space, $(G,\pi)\in\mathbf{PEM}_{V}$, and $W\subset V$ be a subspace. We define the sets
$L_{(G,\pi)}(W)$ :=$:=\{x\in G\colon P_{W}\pi(x)=\pi(x)P_{W}\},$
$S_{(G,\pi)}(W)$ :=$:=\{x\in G\colon P_{W}\pi(x)P_{W}\in\mathbb{T}P_{W}\},$
where $P_{W}$ denotes the orthogonal projection onto $W$. One can of course make analogous definitions for error models in place of projective error models.
Notice that both $L_{(G,\pi)}(W)$ and $S_{(G,\pi)}(W)$ are subgroups of $G$, and in fact $S_{(G,\pi)}(W)$ is a subgroup of $L_{(G,\pi)}(W)$. The group $L_{(G,\pi)}(W)$ is the group of logical operations on the code $W$, and the group $S_{(G,\pi)}(W)$ is the group of stabilizers.
We now want to consider a class of codes that are nice enough so that we may describe the set $D_{(G,\pi)}(W)$ of detectable errors in terms of these two groups.
Proposition 5.2.
Let $V$ be a Hilbert space, $(G,\pi)\in\mathbf{PEM}_{V}$, and $W\subset V$ be a subspace. Then the following conditions are equivalent:
(3) $\pi(x)W\subset W\text{ or }\pi(x)W\subset W^{\perp}\text{ for every }x\in G.$
(4) $D_{(G,\pi)}(W)=(G\setminus L_{(G,\pi)}(W))\cup S_{(G,\pi)}(W).$
Proof.
Suppose that $\pi(x)W\subset W$ or $\pi(x)W\subset W^{\perp}$ for every $x\in G$. If $x\in G$ is such that $\pi(x)W\subset W$, then $x\in L_{(G,\pi)}(W)$, hence $x\in D_{(G,\pi)}(W)$ if and only if $x\in S_{(G,\pi)}(W)$. Furthermore, if $x\in G\setminus L_{(G,\pi)}(W)$ then $\pi(x)W\subset W^{\perp}$, and it is clear that these are precisely the elements of $D_{(G,\pi)}(W)$ that satisfy $P_{W}\pi(x)P_{W}=0$.
For the converse, it is clear that if $x\in L_{(G,\pi)}(W)$, then $\pi(x)W\subset W$. If $x\in G\setminus L_{(G,\pi)}(W)$, then $x\in D_{(G,\pi)}(W)$ and satisfies $P_{W}\pi(x)P_{W}=0$. From this it is clear that $\pi(x)W\subset W^{\perp}$, completing the proof. ∎
This result should really be thought of as giving a condition ensuring that the detectable errors of a code can be described in terms of the stabilizers and logical operations on the code.
Having Equation 4 in mind, we get a new way of thinking about stabilizer codes and Clifford codes. A Clifford code is a code $W$ where one has prescribed the group $L_{(G,\pi)}(W)$ of logical operations, while a stabilizer code is a code $W$ where one has prescribed a subgroup of stabilizers $S_{(G,\pi)}(W)$. We make this precise in the following Theorem, which recovers [KlappeneckerRöttelerII02, Theorem 1] adapted to the present setting.
Theorem 5.3.
Let $V$ be a Hilbert space and $(G,\pi)\in\mathbf{PEM}_{V}$. Suppose that $W=V^{Cl}(G,\pi,I,\rho)$ is a Clifford code. Then
$L_{(G,\pi)}(W)$ $=I,$
$S_{(G,\pi)}(W)$ $=\ker(q\circ\rho),$
$D_{(G,\pi)}(W)$ $=(G\setminus I)\cup\ker(q\circ\rho).$
Proof.
Note first that by the definition of a Clifford code, we have that $I\subset L_{(G,\pi)}(W)$. To see that they are equal we will show that for any $x\in G\setminus I$, $P_{W}\pi(x)P_{W}=0$.
Let $T_{1}\colon V\to\operatorname{Ind}^{G}_{I}V_{\rho}$ be an intertwiner exhibiting the isomorphism $\pi\simeq\operatorname{Ind}^{G}_{I}\rho$, and similarly, let $T_{2}\colon W\to V_{\rho}$ be an intertwiner exhibiting the isomorphism $\operatorname{Res}^{G}_{I}\pi|_{W}\simeq\rho$. Then we have a commuting square of intertwiners:
[…]
Here $i\colon W\to V$ is the inclusion, and $j\colon V_{\rho}\to\operatorname{Ind}^{G}_{I}V_{\rho}$ is the map defined by $j(\eta)=1_{G}\otimes\eta$. Let $x\in G$ and write $x=ry$ with $y\in I$ and $r\in G\setminus I$. If $\xi\in W$, then we compute that
$P_{W}\pi(x)i(\xi)$ $=P_{W}\pi(x)T_{1}^{-1}T_{1}(i(\xi))$
$=P_{W}T_{1}^{-1}\operatorname{Ind}^{G}_{I}\rho(x)T_{1}(i(\xi))$
$=P_{W}T_{1}^{-1}\operatorname{Ind}^{G}_{I}\rho(x)j(T_{2}(\xi))$
$=P_{W}T_{1}^{-1}\operatorname{Ind}^{G}_{I}\rho(x)(1_{G}\otimes T_{2}(\xi))$
$=P_{W}T_{1}^{-1}(x\otimes T_{2}(\xi))$
$=P_{W}T_{1}^{-1}(ry\otimes T_{2}(\xi))$
$=P_{W}T_{1}^{-1}(r\otimes\rho(y)T_{2}(\xi)).$
Notice that $P_{W}T_{1}^{-1}(r\otimes\rho(y)T_{2}(\xi))$ is zero if and only if $r\otimes\rho(y)T_{2}(\xi)$ is not of the form $j(\eta)$ for any $\eta\in V_{\rho}$. This happens if and only if $r\neq 1_{G}$, which is true if and only if $x\in G\setminus I$. Hence, $P_{W}\pi(x)P_{W}=0$ whenever $x\in G\setminus I$ proving that
$L_{(G,\pi)}(W)=I.$
Furthermore, this shows that $\pi(x)W\subset W$ for every $x\in I$ and that $\pi(x)W\subset W^{\perp}$ for every $x\in G\setminus I$. Hence, by Proposition 5.2 we get that
$D_{(G,\pi)}(W)=(G\setminus L_{(G,\pi)}(W))\cup S_{(G,\pi)}(W).$
It remains to show that $S_{(G,\pi)}(W)=\ker(q\circ\rho)$.
Now suppose that $P_{W}T_{1}^{-1}(r\otimes\rho(y)T_{2}(\xi))$ is a nonzero scalar multiple of $\xi$. By the above argument we know that $x\in I$, hence
$P_{W}\pi(x)i(\xi)=P_{W}iT_{2}^{-1}(\rho(x)T_{2}(\xi)).$
This is a nonzero scalar multiple of $\xi$ if and only if $\rho(x)$ is a scalar multiple of $1_{U(V_{\rho})}$, i.e., if and only if $x\in\ker(q\circ\rho)$. This shows that
$S_{(G,\pi)}(W)=\ker(q\circ\rho),$
completing the proof. ∎
Corollary 5.4.
Let $V$ be a Hilbert space and $(G,\pi)\in\mathbf{PEM}_{V}$. If $W$ is a Clifford code with
$V^{Cl}(G,\pi,I_{1},\rho_{1})=W=V^{Cl}(G,\pi,I_{2},\rho_{2}),$
then $I_{1}=I_{2}$ and $\rho_{1}\simeq\rho_{2}$.
Proof.
By Theorem 5.3 we get that $I_{1}=L_{(G,\pi)}(W)=I_{2}$. Furthermore, we also get that $\rho_{1}\simeq(\operatorname{Res}^{G}_{I_{1}}\pi)|_{W}\simeq\rho_{2}$, completing the proof. ∎
Corollary 5.5.
Let $V$ be a Hilbert space and $(G,\pi)\in\mathbf{PEM}_{V}$. If $W$ is a Clifford code then we have that
$|L_{(G,\pi)}(W)|=\frac{\dim W}{\dim V}|G|.$
Proof.
By Theorem 5.3 we have that $\operatorname{Ind}^{G}_{L_{(G,\pi)}(W)}((\operatorname{Res}^{G}_{L_{(G,\pi)}(W)}\pi)|_{W})\simeq\pi$. This implies that
$\dim V=\dim(\operatorname{Ind}^{G}_{L_{(G,\pi)}(W)}W)=\frac{|G|}{|L_{(G,\pi)}(W)|}\dim W.$
Rearranging this equation completes the proof. ∎
With the subgroup $S_{(G,\pi)}(W)$ at hand we can state some more properties of weak stabilizer codes that will turn out to be quite useful.
Lemma 5.6.
Let $V$ be a Hilbert space and $(G,\pi)\in\mathbf{PEM}_{V}$. If $W=V^{St}(G,\pi,H,f)$ is a weak stabilizer code, then $H\subset S_{(G,\pi)}(W)$.
Proof.
For every $x\in H$ we have that $P_{W}\pi(x)P_{W}=f(x)P_{W}$, hence $x\in S_{(G,\pi)}(W)$. ∎
Proposition 5.7.
Let $V$ be a Hilbert space and $(G,\pi)\in\mathbf{PEM}_{V}$. If $W=V^{St}(G,\pi,H,f)$ is a weak stabilizer code, then there exists a function $\tilde{f}\colon S_{(G,\pi)}(W)\to\mathbb{T}$ such that $W=V^{St}(G,\pi,S_{(G,\pi)}(W),\tilde{f})$.
Proof.
For $x\in S_{(G,\pi)}(W)$ define $\tilde{f}(x)$ to be the unique scalar such that $P_{W}\pi(x)P_{W}=\tilde{f}(x)P_{W}$. By Lemma 5.6 we have that $H\subset S_{(G,\pi)}(W)$, hence
$V^{St}(G,\pi,S_{(G,\pi)}(W),\tilde{f})\subset V^{St}(G,\pi,H,f)=W.$
On the other hand, by the definition of $S_{(G,\pi)}(W)$, we have that
$W\subset V^{St}(G,\pi,S_{(G,\pi)}(W),\tilde{f}),$
hence they are equal, completing the proof. ∎
Theorem 5.3 can be used to determine if a weak stabilizer code is also a Clifford code, as this result gives strict criteria on the form of the groups $L_{(G,\pi)}(W)$ and $S_{(G,\pi)}(W)$. Similarly, Proposition 5.7 is helpful to determine if a Clifford code is also a weak stabilizer code, as a Clifford code $W=V^{Cl}(G,\pi,I,\rho)$ need not be equal to $V^{St}(G,\pi,S_{(G,\pi)}(W),\tilde{f})$.
Furthermore, Lemma 5.6 gives an answer to the second of our original questions: given some noise we want to protect against, what is the largest code space that can detect this given noise? If $W=V^{St}(G,\pi,H,f)$ is a weak stabilizer code, then $H\subset S_{(G,\pi)}(W)\subset D_{(G,\pi)}(W)$, hence the code space $W$ can detect all the noise that is being modeled by the subgroup $H$ of $G$. It is also clear from the definition of a weak stabilizer code that $W$ is the largest code space that can detect this noise.
6. Weak stabilizer Clifford codes
In this section we give a complete characterization of when a Clifford code is also a weak stabilizer code for projective error models $(G,\pi)\in\mathbf{PEM}_{V}$ where $|G|=(\dim V)^{2}$, i.e., $G$ is a group of central type. This generalizes [KlappeneckerRötteler04, Theorem 3] and [GrasslMartínezNicolás10, Corollary 4.2]. We need the following elementary fact about the characters of groups of central type. This fact can be established from results found within [CeccheriniSilbersteinTullio22] and [Cheng15], for the ease of the reader we choose to present a proof.
Lemma 6.1.
Let $G$ be a group of central type, and $\pi\colon G\to U(V)$ be an irreducible projective representation such that $|G|=(\dim V)^{2}$. Then we have that $\chi_{\pi}(x)=0$ if and only if $x\neq 1_{G}$.
Proof.
By Schur’s Lemma and Equation 2 we have that
11 11 $=\dim\operatorname{Hom}_{G}(\pi,\pi)$
$=\langle\chi_{\pi},\chi_{\pi}\rangle$
$=\frac{1}{|G|}\sum_{x\in G}\overline{\chi_{\pi}(x)}\chi_{\pi}(x)$
$=\frac{\overline{\chi_{\pi}(1_{G})}\chi_{\pi}(1_{G})}{|G|}+\frac{1}{|G|}\sum_{x\in G\setminus\{1_{G}\}}\overline{\chi_{\pi}(x)}\chi_{\pi}(x)$
$=\frac{(\dim V)^{2}}{|G|}+\frac{1}{|G|}\sum_{x\in G\setminus\{1_{G}\}}\overline{\chi_{\pi}(x)}\chi_{\pi}(x)$
$=1+\frac{1}{|G|}\sum_{x\in G\setminus\{1_{G}\}}\overline{\chi_{\pi}(x)}\chi_{\pi}(x).$
Hence, $\chi_{\pi}(x)=0$ for all $x\in G\setminus\{1_{G}\}$, completing the proof. ∎
Theorem 6.2.
Let $V$ be a Hilbert space and $(G,\pi)\in\mathbf{PEM}_{V}$ with $|G|=(\dim V)^{2}$. If $W\subset V$ is a Clifford code, then $W$ is a weak stabilizer code if and only if $|G|=|L_{(G,\pi)}(W)|\cdot|S_{(G,\pi)}(W)|$.
Proof.
Suppose that $W$ is a weak stabilizer code. Let $H$ be a subgroup of $G$ and $f\colon H\to\mathbb{T}$ be a function such that $W=V^{St}(G,\pi,H,f)$. By Corollary 5.5 we have that
$\dim W=\frac{|L_{(G,\pi)}(W)|}{|G|}\dim V.$
By Proposition 4.7, Equation 2, and Lemma 6.1 we also have that
$\dim W$ $=\dim\operatorname{Hom}_{H}(f,\operatorname{Res}^{G}_{H}\pi)$
$=\langle f,\chi_{\operatorname{Res}^{G}_{H}\pi}\rangle$
$=\frac{1}{|H|}\sum_{x\in H}\overline{f(x)}\chi_{\pi}(x)$
$=\frac{1}{|H|}\dim V.$
By comparing these two equations for $\dim W$ we get that $|G|=|L_{(G,\pi)}(W)|\cdot|H|$. Since $H$ was any subgroup of $G$ displaying the fact that $W$ is a weak stabilizer code, we get that this equality holds for all subgroups of $G$ that exhibits $W$ as a weak stabilizer code. In particular, by Lemma 5.6 we have that $H\subset S_{(G,\pi)}(W)$, and by Proposition 5.7 we have that $S_{(G,\pi)}(W)$ is a subgroup of $G$ that exhibits $W$ as a weak stabilizer code. Hence, $|H|=|S_{(G,\pi)}(W)|$ and in fact $H=S_{(G,\pi)}(W)$.
To prove the converse statement, suppose that $|G|=|L_{(G,\pi)}(W)|\cdot|S_{(G,\pi)}(W)|$. By definition of $S_{(G,\pi)}(W)$ there exist a function $f\colon S_{(G,\pi)}(W)\to\mathbb{T}$ such that
$W\subset W^{\prime}:=V^{St}(G,\pi,S_{(G,\pi)}(W),f).$
By the same computations as above we have that
$\dim W=\frac{|L_{(G,\pi)}(W)|}{|G|}\dim V=\frac{1}{|S_{(G,\pi)}(W)|}\dim V=\dim W^{\prime}.$
Hence, $W=W^{\prime}$, completing the proof. ∎
We choose to highlight the following facts which were shown in the proof above.
Proposition 6.3.
Let $V$ be a Hilbert space and $(G,\pi)\in\mathbf{PEM}_{V}$ with $|G|=(\dim V)^{2}$. If $W\subset V$ is a Clifford code and a weak stabilizer code, and $W=V^{St}(G,\pi,H,f)$ for some function $f\colon H\to\mathbb{T}$, then $H=S_{(G,\pi)}(W)$.
Proposition 6.4.
Let $V$ be a Hilbert space and $(G,\pi)\in\mathbf{PEM}_{V}$ with $|G|=(\dim V)^{2}$. If $W\subset V$ is a non-stabilizer Clifford code, then
$|L_{(G,\pi)}(W)|\cdot|S_{(G,\pi)}(W)|<|G|.$
From this we immediately get the following result.
Corollary 6.5.
Let $V$ be a Hilbert space and $(G,\pi)\in\mathbf{PEM}_{V}$ with $|G|=(\dim V)^{2}$. If $W\subset V$ is a Clifford code, then $W$ is a stabilizer code if and only if $S_{(G,\pi)}(W)$ is a normal subgroup of $G$ and $|G|=|L_{(G,\pi)}(W)|\cdot|S_{(G,\pi)}(W)|$.
Proof.
By Theorem 6.2 we have that $W$ is a weak stabilizer code if and only if
$|G|=|L_{(G,\pi)}(W)|\cdot|S_{(G,\pi)}(W)|.$
In this case, we have by Proposition 6.3 that $S_{(G,\pi)}(W)$ is the only subgroup of $G$ for which there exists a function $f\colon S_{(G,\pi)}(W)\to\mathbb{T}$ such that
$W=V^{St}(G,\pi,S_{(G,\pi)}(W),f).$
Hence, $W$ is a stabilizer code if and only if $S_{(G,\pi)}(W)$ is a normal subgroup of $G$. This completes the proof. ∎
7. Examples of projective error models
In this section we will give some important examples of some of the concepts we have talked about so far. First we would like to point out that for any Hilbert space $V$, the collection of projective error models $\mathbf{PEM}_{V}$ is non-empty. This is of course well known, but it is nice to have it highlighted.
Proposition 7.1.
Let $n\geq 1$ be an integer, $\zeta_{n}\in\mathbb{C}$ be a primitive $n$-th root of unity, and consider the group $\mathbb{Z}_{n}\times\mathbb{Z}_{n}$. Then the function $\pi\colon\mathbb{Z}_{n}\times\mathbb{Z}_{n}\to U(\mathbb{C}^{n})$ defined by
$\pi(a,b)=X_{n}^{a}Z_{n}^{b},$
where
$X_{n}=\begin{bmatrix}0&1&0&\cdots&0&0\\ 0&0&1&\cdots&0&0\\ 0&0&0&\cdots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\cdots&0&1\\ 1&0&0&\cdots&0&0\end{bmatrix},\text{ and }Z_{n}=\begin{bmatrix}1&0&0&\cdots&0&0\\ 0&\zeta_{n}&0&\cdots&0&0\\ 0&0&\zeta_{n}^{2}&\cdots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\cdots&\zeta_{n}^{n-2}&0\\ 0&0&0&\cdots&0&\zeta_{n}^{n-1}\end{bmatrix},$
is a projectively faithful irreducible projective representation of $\mathbb{Z}_{n}\times\mathbb{Z}_{n}$ on $\mathbb{C}^{n}$.
Proof.
One computes that $(X_{n})^{n}=(Z_{n})^{n}=I_{n}$ and that $X_{n}Z_{n}=\zeta_{n}Z_{n}X_{n}$. Hence, the defining relations for $\mathbb{Z}_{n}\times\mathbb{Z}_{n}$ are satisfied by the elements $q(X_{n}),q(Z_{n})\in PU(\mathbb{C}^{n})$. Thus, the composition $q\circ\pi\colon\mathbb{Z}_{n}\times\mathbb{Z}_{n}\to PU(\mathbb{C}^{n})$ is a group homomorphism, meaning that $\pi$ is a projective representation of $\mathbb{Z}_{n}\times\mathbb{Z}_{n}$ on $\mathbb{C}^{n}$.
Notice that the set $\{\pi(a,b)\colon a,b\in\mathbb{Z}_{n}\}$ is linearly independent, so since the cardinality of $\mathbb{Z}_{n}\times\mathbb{Z}_{n}$ is $n^{2}$ we get that this is in fact a basis for $M_{n}(\mathbb{C})$. This shows that $\pi$ is projectively faithful and irreducible, completing the proof. ∎
The next class of projective error models is derived from the XP-stabilizer formalism, cf. [Webster2022]. We would also like to mention that the following facts about the dihedral groups are well known, and an exposition can be found in [Karpilovsky85, Chapter 3.7]. For the ease of the reader we will present a proof here.
Proposition 7.2.
Let $n\geq 2$ be an integer, $\zeta_{n}\in\mathbb{C}$ be a primitive $n$-th root of unity, and consider the dihedral group of $2n$ elements with the following presentation:
$D_{n}=\langle a,b\,|\,a^{n}=b^{2}=1,\,bab=a^{n-1}\rangle=\{1,a,a^{2},\dots,a^{n-1},b,ba,\dots,ba^{n-1}\}.$
Then the function $\pi\colon D_{n}\to U(\mathbb{C}^{2})$ defined by
$\pi(b^{k}a^{l})=X^{k}P^{l},\text{ for }k=0,1\text{ and }l=0,1,\dots,n-1,$
where
$X=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},\text{ and }P=\begin{bmatrix}1&0\\ 0&\zeta_{n}\end{bmatrix},$
is a projectively faithful irreducible projective representation of $D_{n}$ on $\mathbb{C}^{2}$.
Proof.
We first compute that $P^{n}=X^{2}=I$ and that $XPX=\zeta_{n}P^{n-1}$. This means that the defining relations for $D_{n}$ are satisfied by the elements $q(P),q(X)\in PU(\mathbb{C}^{2})$. Thus, the composition $q\circ\pi\colon D_{n}\to PU(\mathbb{C}^{2})$ is a group homomorphism, meaning that $\pi$ is a projective representation of $D_{n}$ on $\mathbb{C}^{2}$.
Notice further that the elements $I,X,P,XP$ are linearly independent, hence they form a basis for $M_{2}(\mathbb{C})$. In particular this means that $\operatorname{span}\{\pi(x)\colon x\in D_{n}\}=M_{2}(\mathbb{C})$, hence $\pi$ is irreducible.
Finally, one easily computes that $\ker(q\circ\pi)=\{1\}$, hence $\pi$ is projectively faithful, completing the proof. ∎
Remark 7.3.
One can compute that the 2-cocycle associated to the above projective representation is given by the function $\sigma\colon D_{n}\times D_{n}\to\mathbb{T}$, defined by
$\sigma(b^{k_{1}}a^{l_{1}},b^{k_{2}}a^{l_{2}})=\zeta_{n}^{k_{2}l_{1}}.$
Furthermore, the second cohomology group
$H^{2}(D_{n},\mathbb{T})=\begin{cases}0,&n\text{ is odd},\\ \mathbb{Z}_{2},&n\text{ is even}.\end{cases}$
Hence, when $n$ is odd, $\sigma$ is a coboundary. In fact, if we define $f\colon D_{n}\to\mathbb{T}$ by
$f(b^{k}a^{l})=\zeta_{n}^{(n-1)l/2},$
then $\sigma\,(\delta f)=1$.
Given two projective error models, there is a natural way of combining them to give us a new projective error model. In what follows we explain how to construct such product projective error models.
Proposition 7.4.
Let $V_{1}$ and $V_{2}$ be Hilbert spaces and suppose that $(G_{1},\pi_{1})\in\mathbf{PEM}_{V_{1}}$ and $(G_{2},\pi_{2})\in\mathbf{PEM}_{V_{2}}$. Then the projective representation $\pi_{1}\otimes\pi_{2}\colon G_{1}\times G_{2}\to U(V_{1}\otimes V_{2})$ defined by
$(\pi_{1}\otimes\pi_{2})(x,y)=\pi_{1}(x)\otimes\pi_{2}(y),\ x\in G_{1},\text{ and }y\in G_{2},$
is a projectively faithful irreducible projective representation of $G_{1}\times G_{2}$ on $V_{1}\otimes V_{2}$.
Proof.
It is straightforward to check that the projective representation $\pi_{1}\otimes\pi_{2}$ is irreducible. One can for example confirm that $\operatorname{span}\{(\pi_{1}\otimes\pi_{2})(x,y)\colon x\in G_{1},\ y\in G_{2}\}=B(V_{1}\otimes V_{2})$.
To see that $\pi_{1}\otimes\pi_{2}$ is faithful, suppose that $(\pi_{1}\otimes\pi_{2})(x,y)$ is a scalar multiple of the identity. Then both $\pi_{1}(x)$ and $\pi_{2}(y)$ are scalar multiples of the identities on $V_{1}$ and $V_{2}$ respectively. Since both $\pi_{1}$ and $\pi_{2}$ are assumed to be projectively faithful we can conclude that $x=1_{G_{1}}$ and $y=1_{G_{2}}$. This completes the proof. ∎
This fact is also well known; see for example [Serre77, Theorem 10] for a proof in the linear setting.
With this established we see that the projective Pauli error model, as explained in Example 3.6, arise by considering either of the projective error models in Proposition 7.1 or Proposition 7.2 with $n=2$, and taking products of these models with themselves as explained in Proposition 7.4.
When taking products as above we can also incorporate an action from the symmetric group on $n$ letters, $S_{n}$. This gives a more sophisticated way of constructing projective error models, as has been done in for example [Pollatsek-Ruskai04] and [AydinAlekseyevBarg24] where the authors construct permutation invariant codes that can correct errors from the Pauli group.
Proposition 7.5.
Let $V$ be a Hilbert space and suppose that $(G,\pi)\in\mathbf{PEM}_{V}$. For $n\geq 1$ let $S_{n}$ act on $G^{n}$ by permuting the $n$ copies of $G$. Then the semi direct product $G^{n}\rtimes S_{n}$ constructed from this action admits a projectively faithful irreducible projective representation on $V^{\otimes n}$.
Proof.
There is an obvious faithful representation $S_{n}\to U(V^{\otimes n})$ given by permuting the tensor factors. By abuse of notation we will just denote this representation by $\tau\mapsto\tau$, for any $\tau\in S_{n}$. The function $\pi^{n}_{Sym}\colon G^{n}\rtimes S_{n}\to U(V^{\otimes n})$ defined by
$\pi^{n}_{Sym}(x_{1},\dots,x_{n},\tau)=(\pi(x_{1})\otimes\dots\otimes\pi(x_{n}))\,\tau$
is then easily verified to be a projective representation of $G^{n}\rtimes S_{n}$ on $V^{\otimes n}$. Furthermore, it is faithful if and only if $\pi$ is faithful, and it is irreducible if $\pi$ is irreducible, completing the proof. ∎
8. Examples of non-stabilizer Clifford codes
In [KlappeneckerRötteler02, Section 10.9] the first and smallest example of a non-stabilizer Clifford code is constructed. Since the authors mainly were concerned with projective error models from groups of central type, or nice error bases, it is not mentioned that this example fits into an infinite family of similar examples. Inserting $n=2$ in the next example retrieves the well known example by Klappenecker and Rötteler.
Proposition 8.1.
Let $n\geq 2$ be an integer and consider the group $C_{2}\times D_{2n}$ with the following presentation:
$C_{2}\times D_{2n}=\langle a,b,c\,|\,a^{2n}=b^{2}=c^{2}=[a,c]=[b,c]=1,\,bab=a^{-1}\rangle.$
Let $X$ and $P$ be the matrices as in Proposition 7.2, and $\rho\colon D_{2n}\to U(\mathbb{C}^{2})$ be the corresponding projective representation. Define $\pi\colon C_{2}\times D_{2n}\to U(\mathbb{C}^{4})$ by
$\pi(c^{k}b^{l}a^{m})=\begin{bmatrix}0&I_{2}\\ I_{2}&0\end{bmatrix}^{k}\begin{bmatrix}X&0\\ 0&X\end{bmatrix}^{l}\begin{bmatrix}P&0\\ 0&-P\end{bmatrix}^{m}.$
Then $\pi$ is a projectively faithful irreducible projective representation of $C_{2}\times D_{2n}$ on $\mathbb{C}^{4}$. Moreover, $\operatorname{Ind}_{D_{2n}}^{C_{2}\times D_{2n}}\rho\simeq\pi$, and $W:=V^{Cl}(C_{2}\times D_{2n},\pi,D_{2n},\rho)$ is a non-stabilizer Clifford code.
Proof.
As in the proof of Proposition 7.2 one computes that $\pi$ is indeed a projective representation of $C_{2}\times D_{2n}$ on $\mathbb{C}^{4}$. To see that $\pi$ is irreducible one confirms that
$\operatorname{span}\{\pi(x)\colon x\in C_{2}\times D_{2n}\}=M_{4}(\mathbb{C}).$
To see that $\pi$ is projectively faithful, notice that $\pi(c^{k}b^{l}a^{m})$ is a diagonal matrix if and only if $k=l=0$. In this case it is easy to see that $\pi(c^{k}b^{l}a^{m})$ is a scalar multiple of the identity if and only if $m=0$ as well, hence $\pi$ is projectively faithful.
To see that $\operatorname{Ind}_{D_{2n}}^{C_{2}\times D_{2n}}\rho\simeq\pi$ define the linear map $T\colon\operatorname{Ind}_{D_{2n}}^{C_{2}\times D_{2n}}\mathbb{C}^{2}\to\mathbb{C}^{4}$, by
$T(x\otimes(z_{1},z_{2}))=\pi(x)(z_{1},z_{2},0,0).$
This is an intertwiner, and we notice that the Clifford code $W$ corresponds to the first two coordinates in $\mathbb{C}^{4}$.
By Proposition 7.2 we know that $\rho$ is projectively faithful. Hence, by Theorem 5.3 we get that
$S_{(C_{2}\times D_{2n},\pi)}(W)=\ker(q\circ\rho)=\{1\}.$
Then
$V^{St}(C_{2}\times D_{2n},\pi,S_{(C_{2}\times D_{2n},\pi)}(W),1)=\mathbb{C}^{4}.$
By Proposition 5.7 this implies that $W$ is not a weak stabilizer code, completing the proof. ∎
The next family of examples is inspired by the above example by Klappenecker and Rötteler, as well as the work by Ginosar and Schnabel in [GinosarSchnabel19, Theorem A], where they present a large family of groups of central type.
Proposition 8.2.
Let $n\geq 3$ be an odd number. Let $\mathbb{Z}_{2}$ act on $\mathbb{Z}_{n}\times\mathbb{Z}_{n}$ via inversion, and let $I:=(\mathbb{Z}_{n}\times\mathbb{Z}_{n})\rtimes\mathbb{Z}_{2}$ denote the associated semi-direct product. Define $\rho\colon I\to U(\mathbb{C}^{n})$ by
$\rho(a,b,c)=X_{n}^{a}Z_{n}^{b}C^{c},$
where $X_{n}$ and $Z_{n}$ are as in Proposition 7.1 and
$C:=\begin{bmatrix}0&0&\cdots&0&1\\ 0&0&\cdots&1&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&1&\cdots&0&0\\ 1&0&\cdots&0&0\end{bmatrix}.$
Set $G:=I\times\mathbb{Z}_{2}$ and define $\pi\colon G\to U(\mathbb{C}^{2n})$ by
$\pi(a,b,c,d)=\begin{bmatrix}X_{n}&0\\ 0&X_{n}\end{bmatrix}^{a}\begin{bmatrix}Z_{n}&0\\ 0&Z_{n}\end{bmatrix}^{b}\begin{bmatrix}C&0\\ 0&-C\end{bmatrix}^{c}\begin{bmatrix}0&I_{n}\\ I_{n}&0\end{bmatrix}^{d}.$
Then both $\rho$ and $\pi$ are projectively faithful irreducible projective representations, and $\operatorname{Ind}_{I}^{G}\rho\simeq\pi$. Furthermore, $W:=V^{Cl}(G,\pi,I,\rho)$ is a non-stabilizer Clifford code.
Proof.
The fact that $\rho$ is a projectively faithful irreducible projective representation of $I$ follows by Proposition 7.1. To see that $\pi$ is a projectively faithful irreducible projective representations of $G$, we claim that the collection $\{\pi(x):x\in G\}$ forms a basis for $M_{2n}(\mathbb{C})$. Indeed, if $\pi(x)$ and $\pi(y)$ are linearly dependent, then we get that $\pi(x^{-1}y)$ is a scalar multiple of the identity matrix. Notice that $\pi(a,b,c,d)$ is a diagonal matrix if and only if $a=c=d=0$. From this observation it becomes clear that $\pi(x^{-1}y)$ is a scalar multiple of the identity if and only if $x^{-1}y=1_{G}$. Hence, $x=y$ and thus $\pi(x)=\pi(y)$. This shows that the set $\{\pi(x):x\in G\}$ is linearly independent, hence a basis for $M_{2n}(\mathbb{C})$ since $|G|=(2n)^{2}$.
The fact that $\operatorname{Ind}_{I}^{G}\rho\simeq\pi$ follows by essentially the same reasoning as in Proposition 8.1, so does the fact that $W=V^{Cl}(G,\pi,I,\rho)$ is a non-stabilizer Clifford code. This completes the proof. ∎
9. Examples of non-Clifford weak stabilizer codes
To find examples of non-Clifford weak stabilizer codes we want to use Theorem 5.3 to its full potential. The following example is an example of a code that is permutation invariant. We thank Ningping Cao for the idea that lead us towards this example.
Proposition 9.1.
Let $n\geq 2$ and define $G=(\mathbb{Z}_{2}\times\mathbb{Z}_{2})^{n}\rtimes S_{n}$ where $S_{n}$ acts on $(\mathbb{Z}_{2}\times\mathbb{Z}_{2})^{n}$ by permuting the copies of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. Consider the projectively faithful irreducible projective representation $\pi\colon G\to U((\mathbb{C}^{2})^{\otimes n})$ defined by
$\pi((a_{i},b_{i})_{i=1}^{n},\tau)=(X^{a_{1}}Z^{b_{1}}\otimes\dots\otimes X^{a_{n}}Z^{b_{n}})\,\tau,$
where $X$ and $Z$ are defined in Example 3.6. Then $W:=V^{St}(G,\pi,S_{n},1)$ is a weak stabilizer code that is not a Clifford code.
Proof.
Note that $\pi$ is indeed a projectively faithful irreducible projective representation by Proposition 7.5. Let $\left|0\right\rangle=(1,0),\left|1\right\rangle=(0,1)$ denote the standard basis vectors on $\mathbb{C}^{2}$. We will use the shorthand notation
$\left|j_{1}j_{2}\dots j_{n}\right\rangle:=\left|j_{1}\right\rangle\otimes\left|j_{2}\right\rangle\otimes\dots\otimes\left|j_{n}\right\rangle,$
with each $j_{k}\in\{0,1\}$, to denote the standard basis for $(\mathbb{C}^{2})^{\otimes n}$. For each $k=0,1,\dots,n$ we define the set
$B_{k}=\{\left|j_{1}j_{2}\dots j_{n}\right\rangle\colon\sum_{i=1}^{n}j_{i}=k\},$
and define the vector
$\left|D^{n}_{k}\right\rangle=\sum_{\xi\in B_{k}}\xi.$
One easily confirms that $W=\operatorname{span}\{\left|D^{n}_{k}\right\rangle\colon k=0,1,\dots,n\}$, hence $W$ is in fact a weak stabilizer code as it is a nonzero subspace. If $W$ is a Clifford code, we have by Corollary 5.5 that
$|L_{(G,\pi)}(W)|=\frac{\dim W}{\dim((\mathbb{C}^{2})^{\otimes n})}|G|=2^{n}\cdot(n+1)!$
Suppose that $((a_{i},b_{i})_{i=1}^{n},\tau)\in L_{(G,\pi)}(W)$. Then we have that
$\pi((a_{i},b_{i})_{i=1}^{n},\tau)\left|D_{0}^{n}\right\rangle=\left|a_{1}a_{2}\dots a_{n}\right\rangle$
must belong to $W$. However, this is the case if and only if $a_{1}=a_{2}=\dots=a_{n}$. Thus, we get that
$|L_{(G,\pi)}(W)|\leq 2^{n+1}\cdot n!<2^{n}\cdot(n+1)!$
whenever $12<n+$ 1. Hence, whenever $n\geq 2$, the order of $L_{(G,\pi)}(W)$ is too small for $W$ to be a Clifford code, completing the proof. ∎
These examples of non-Clifford weak stabilizer codes are very badly behaved. For example, when $n=2$ the vector
$(X\otimes I_{2})\left|D^{2}_{0}\right\rangle=\left|10\right\rangle$
is not in the code space or orthogonal to it. Hence, the assumptions in Proposition 5.2 are not satisfied, so we cannot describe the detectable errors in terms of the stabilizers and logical operations on the code.
10. Product codes
In this section we show that the tensor product of two Clifford codes is again a Clifford code. In the special case where we only consider projective error models $(G,\pi)\in\mathbf{PEM}_{V}$ where $|G|=(\dim V)^{2}$, we show that if either of the Clifford codes we started with was not a weak stabilizer code, then the product code is also not a weak stabilizer code.
Proposition 10.1.
Let $V_{1}$ and $V_{2}$ be Hilbert spaces, and suppose that $(G_{1},\pi_{1})\in\mathbf{PEM}_{V_{1}}$, and $(G_{2},\pi_{2})\in\mathbf{PEM}_{V_{2}}$. If $W_{1}\subset V_{1}$, and $W_{2}\subset V_{2}$ are Clifford codes, then $W_{1}\otimes W_{2}$ is a Clifford code in $V_{1}\otimes V_{2}$. Furthermore,
$L_{(G_{1}\times G_{2},\pi_{1}\otimes\pi_{2})}(W_{1}\otimes W_{2})$ $=L_{(G_{1},\pi_{1})}(W_{1})\times L_{(G_{2},\pi_{2})}(W_{2}),\text{ and }$
$S_{(G_{1}\times G_{2},\pi_{1}\otimes\pi_{2})}(W_{1}\otimes W_{2})$ $=S_{(G_{1},\pi_{1})}(W_{1})\times S_{(G_{2},\pi_{2})}(W_{2}).$
Proof.
By Proposition 7.4 we have that $(G_{1}\times G_{2},\pi_{1}\otimes\pi_{2})\in\mathbf{PEM}_{V_{1}\otimes V_{2}}$, hence the above statements are well-defined. Write $W_{1}=V^{Cl}(G_{1},\pi_{1},I_{1},\rho_{1})$ and $W_{2}=V^{Cl}(G_{2},\pi_{2},I_{2},\rho_{2})$. To confirm that $W_{1}\otimes W_{2}$ is a Clifford code we want to check that
$\operatorname{Ind}_{I_{1}\times I_{2}}^{G_{1}\times G_{2}}(\rho_{1}\otimes\rho_{2})\simeq(\operatorname{Ind}_{I_{1}}^{G_{1}}\rho_{1})\otimes(\operatorname{Ind}_{I_{2}}^{G_{2}}\rho_{2}).$
Since $\operatorname{Ind}_{I_{1}}^{G_{1}}\rho_{1}\simeq\pi_{1}$ and $\operatorname{Ind}_{I_{2}}^{G_{2}}\rho_{2}\simeq\pi_{2}$, this would prove the desired isomorphism. Define $T\colon\operatorname{Ind}_{I_{1}\times I_{2}}^{G_{1}\times G_{2}}(W_{1}\otimes W_{2})\to(\operatorname{Ind}_{I_{1}}^{G_{1}}(W_{1}))\otimes(\operatorname{Ind}_{I_{2}}^{G_{2}}(W_{2}))$ on simple tensors by
$T((x_{1},x_{2})\otimes(\xi_{1}\otimes\xi_{2}))=(x_{1}\otimes\xi_{1})\otimes(x_{2}\otimes\xi_{2}).$
This is clearly an injective intertwiner. Since $T$ is an injective intertwiner between finite dimensional vector spaces of the same dimension, we get that $T$ is an isomorphism. This shows that $W_{1}\otimes W_{2}$ is a Clifford code, and that
$L_{(G_{1}\times G_{2},\pi_{1}\otimes\pi_{2})}(W_{1}\otimes W_{2})=I_{1}\times I_{2}=L_{(G_{1},\pi_{1})}(W_{1})\times L_{(G_{2},\pi_{2})}(W_{2})$
by Theorem 5.3 and Corollary 5.4. Similarly, we get that
$S_{(G_{1}\times G_{2},\pi_{1}\otimes\pi_{2})}(W_{1}\otimes W_{2})=\ker(q\circ(\rho_{1}\otimes\rho_{2}))=S_{(G_{1},\pi_{1})}(W_{1})\times S_{(G_{2},\pi_{2})}(W_{2}),$
since $\rho_{1}(x_{1})\otimes\rho_{2}(x_{2})$ is a scalar multiple of the identity if and only if both $\rho_{1}(x_{1})$ and $\rho_{2}(x_{2})$ are scalar multiples of the identity. This completes the proof. ∎
Corollary 10.2.
Let $V_{1}$ and $V_{2}$ be Hilbert spaces, and suppose that $(G_{1},\pi_{1})\in\mathbf{PEM}_{V_{1}}$, and $(G_{2},\pi_{2})\in\mathbf{PEM}_{V_{2}}$ are such that $|G_{1}|=(\dim V_{1})^{2}$, and $|G_{2}|=(\dim V_{2})^{2}$. If $W_{1}\subset V_{1}$ and $W_{2}\subset V_{2}$ are Clifford codes such that $W_{1}$ is not a weak stabilizer code, then $W_{1}\otimes W_{2}$ is a non-stabilizer Clifford code.
Proof.
By Proposition 10.1 we have that $W_{1}\otimes W_{2}$ is a Clifford code, and by Theorem 6.2 we have that $W_{1}\otimes W_{2}$ is also a weak stabilizer code if and only if
$|G_{1}\times G_{2}|=|L_{(G_{1}\times G_{2},\pi_{1}\otimes\pi_{2})}(W_{1}\otimes W_{2})|\cdot|S_{(G_{1}\times G_{2},\pi_{1}\otimes\pi_{2})}(W_{1}\otimes W_{2})|.$
By Proposition 10.1 we can rewrite this equation as
$|G_{1}|\cdot|G_{2}|=|L_{(G_{1},\pi_{1})}(W_{1})|\cdot|L_{(G_{2},\pi_{2})}(W_{2})|\cdot|S_{(G_{1},\pi_{1})}(W_{1})|\cdot|S_{(G_{2},\pi_{2})}(W_{2})|.$
But, by Theorem 6.2 and Proposition 6.4 we have that
$|L_{(G_{1},\pi_{1})}(W_{1})|\cdot|S_{(G_{1},\pi_{1})}(W_{1})|$ $<|G_{1}|,\text{ and }$
$|L_{(G_{2},\pi_{2})}(W_{2})|\cdot|S_{(G_{2},\pi_{2})}(W_{2})|$ $\leq|G_{2}|.$
Hence, $W_{1}\otimes W_{2}$ is a non-stabilizer Clifford code, completing the proof. ∎
This results allows us to arbitrarily combine the codes from Proposition 8.2 to produce more examples of non-stabilizer Clifford codes.
11. Further questions and remarks
We have spent some effort in discussing when a Clifford code is a weak stabilizer code, and visa versa. Theorem 6.2 gives a description in the case where $(G,\pi)\in\mathbf{PEM}_{V}$ is a projective error model such that $|G|=(\dim V)^{2}$. A natural question is if such a result is possible in the more general case. To be precise, we state the following question.
Question 11.1.
Let $V$ be a Hilbert space and $(G,\pi)\in\mathbf{PEM}_{V}$. If $W\subset V$ is a Clifford code, is it possible to determine if $W$ is also a weak stabilizer code by knowing the orders $|L_{(G,\pi)}(W)|$, $|S_{(G,\pi)}(W)|$, and $|G|$?
It is also natural to wonder if all codes that are a Clifford code and a weak stabilizer code is automatically a stabilizer code. In other words, is
[…]
To state this question more formally we introduce the following notation. Consider a Hilbert space $V$ and a projective error model $(G,\pi)\in\mathbf{PEM}_{V}$. Define
$St_{(G,\pi)}(V)$ =$=\{W\subset V:W\text{ is a stabilizer code with respect to }(G,\pi)\},$
$wSt_{(G,\pi)}(V)$ =$=\{W\subset V:W\text{ is a weak stabilizer code with respect to }(G,\pi)\},$
$Cl_{(G,\pi)}(V)$ =$=\{W\subset V:W\text{ is a Clifford code with respect to }(G,\pi)\}.$
We have been unable to produce an example of a code that is both a Clifford code and a weak stabilizer code but not a stabilizer code. We therefore state the following question.
Question 11.2.
Does there exist a Hilbert space $V$ and a projective error model $(G,\pi)\in\mathbf{PEM}_{V}$ such that
$St_{(G,\pi)}(V)\neq wSt_{(G,\pi)}(V)\cap Cl_{(G,\pi)}(V)?$
In the special case where $|G|=(\dim V)^{2}$, we can restate this question as follows using Theorem 6.2.
Question 11.3.
Does there exist a Hilbert space $V$, a subspace $W\subset V$, and a projective error model $(G,\pi)\in\mathbf{PEM}_{V}$ with $|G|=(\dim V)^{2}$, such that the equation
$|G|=|L_{(G,\pi)}(W)|\cdot|S_{(G,\pi)}(W)|$
holds, but $S_{(G,\pi)}(W)$ is not normal in $G$?
Finally, it is worth stating that there are some aspects of quantum error correction that we have not discussed but would nevertheless fit into this framework. One of these aspects is the notion of code distance. To define code distance, one would need to consider projective error models of the form $(G^{n},\pi^{\otimes n})$, as explained in Proposition 7.4. See [GrasslMartínezNicolás10, Section 5] where they discuss Hamming weight and Clifford codes that can detect errors of Hamming weight $1$.
Another type of code that we could have discussed is subsystem codes, see [KribsLaflammePoulin2005]. If $V$ is a Hilbert space and $W\subset V$ is a subspace such that $W=W_{1}\otimes W_{2}$ for some Hilbert spaces $W_{1}$ and $W_{2}$, we say that $W_{1}$ is a subsystem code. In this case one would be interested in the following set of detectable errors.
$D^{sub}_{(G,\pi)}(W_{1})=\{x\in G\colon P_{W}\pi(x)P_{W}=\operatorname{id}_{W_{1}}\otimes T_{x}\text{ for some }T_{x}\in\mathcal{B}(W_{2})\}.$
Note that if $W=V^{Cl}(G,\pi,I,\rho)$ is a Clifford code, $I=I_{1}\times I_{2}$, and $\rho\simeq\rho_{1}\otimes\rho_{2}$ for some irreducible projective representations $\rho_{i}\colon I_{i}\to U(W_{i})$, $i=1,2$, then $W\simeq W_{1}\otimes W_{2}$. This gives a natural construction of a subsystem code from a Clifford code. With this, one could attempt to prove a result that is analogous to Theorem 5.3.
[…]